### v1

# SceneVerse: Scaling 3D Vision-Language Learning for Grounded Scene Understanding

Abstract
3D vision-language grounding, which focuses on aligning language with the 3D physical environment, stands as a cornerstone in the development of embodied agents. In comparison to recent advancements in the 2D domain, grounding language in 3D scenes faces several significant challenges: (i) the inherent complexity of 3D scenes due to the diverse object configurations, their rich attributes, and intricate relationships; (ii) the scarcity of paired 3D vision-language data to support grounded learning; and (iii) the absence of a unified learning framework to distill knowledge from grounded 3D data. In this work, we aim to address these three major challenges in 3D vision-language by examining the potential of systematically upscaling 3D vision-language learning in indoor environments. We introduce the first million-scale 3D vision-language dataset, SceneVerse, encompassing about $68$K 3D indoor scenes and comprising $2.5$M vision-language pairs derived from both human annotations and our scalable scene-graph-based generation approach. We demonstrate that this scaling allows for a unified pre-training framework, Grounded Pre-training for Scenes (GPS), for 3D vision-language learning. Through extensive experiments, we showcase the effectiveness of GPS by achieving state-of-the-art performance on all existing 3D visual grounding benchmarks. The vast potential of SceneVerse and GPS is unveiled through zero-shot transfer experiments in the challenging 3D vision-language tasks.
Figure 1: Overview of SceneVerse. A million-scale 3D vision-language dataset that comprises over $68$K various 3D indoor scenes and $2.5$M aligned scene-language pairs in the form of scene caption, object caption, and object referral.
$*{}^{*}$ indicates equal contribution.
1 Introduction
The foundation of human cognitive development lies in the grounding of language within the physical world [73, 46, 97]. Recent progress in Large Language Models (LLMs[11, 75], often referred to as “foundation models” [10], has markedly promoted the alignment between vision and language [66, 3, 51] through utilizing billion-scale vision-language datasets [71, 96]. Nonetheless, with these advancements predominantly focusing on the 2D domain, the grounded understanding of 3D physical environments remains in an incipient stage [16, 1, 5]. Recognizing the pivotal role of grounded 3D experiences in shaping human cognition [7, 8] and the delayed research development in this area, there is a compelling need to intensify the exploration into the vision-language learning challenge, specifically in the context of 3D scenes.
Seeking insights from the 2D vision-language (2D-VL) achievements, a major factor to the success was the notable scale-up of paired vision-language data [45, 15, 71]. However, applying these principles directly from 2D to 3D is fraught with challenges. Primarily, 3D data collection heavily relies on the scanning device, making it inherently much more complex and expensive than gathering 2D images. Despite steady efforts to increase the volume of 3D scene data [23, 87, 58, 9], most datasets remain limited to thousands of scenes, substantially lagging behind the scale of existing 2D datasets. This gap is further widened by the inherent complexities of 3D scenes, which feature a multitude of object instances with diverse attributes, varying arrangements, and intricate inter-object relationships. These unique aspects of 3D scenes not only make the accurate description of objects and their relations more challenging but also considerably increase the number of language descriptions required for thorough scene depiction. Consequently, this presents a significant challenge in obtaining a sufficient supply of high-quality paired scene-language data crucial for grounded scene understanding.
To confront these challenges, we propose consolidating current efforts to build up SceneVerse, the first million-scale dataset aimed at advancing 3D vision-language (3D-VL) learning for grounded scene understanding. At the scene level, we unify 3D scene data from existing datasets [23, 78, 9, 67, 40] and supplement the collection with synthetic scenes [95, 27]. This compilation represents the most extensive 3D scene data gathered to date, amounting to $68,406$ scenes for grounding. Additionally, we propose an automated generation pipeline utilizing 3D scene graphs [4, 79] and LLMs to create comprehensive, high-quality scene-language pairs. This refined collection, including $190,836$ human annotated pairs and totaling $2.5$M scene-language pairs, provides detailed and comprehensive portrayals of both object-level and scene-level descriptions within the 3D scene.
We thoroughly investigate the potential offered by the data scale-up in SceneVerse with large-scale pre-training. Specifically, we present a novel and unified pre-training framework, Grounded Pre-training for Scenes (GPS), which is designed with scene-level and object-level alignment objectives and devoid of auxiliary losses and designs. Through multi-level contrastive alignment, we observe significant performance improvements across all existing 3D visual grounding benchmarks, achieving new state-of-the-art results through a simple and effective pre-training process. Additionally, we unveil the vast possibilities offered by SceneVerse and GPS in 3D-VL tasks in a zero-shot transfer setting. At last, we provide a more comprehensive understanding of the data-scaling effects in SceneVerse through extensive ablative experiments to point out future directions.
Our main contributions can be summarized as follows:
  • We introduce SceneVerse, the first million-scale 3D-VL dataset for grounded scene understanding. SceneVerse encompasses $68$K 3D scenes coupled with $2.5$M scene-language pairs, sourced through a combination of human annotation and automated generation methods. This represents a significant improvement in terms of data diversity and scale compared to prior datasets.
  • We propose GPS, an efficient transformer-based model trained with multi-level scene-text alignment that achieves state-of-the-art results on all existing 3D-VL grounding benchmarks, benefiting from pre-training on multi-level scene-language pairs in SceneVerse.
  • We demonstrate that with the data scale-up and model design, our pre-trained models exhibit emerging zero-shot generalization capabilities in grounded scene understanding, paralleling the successes seen in 2D-VL models.
2 Related Work
Datasets for Grounded 3D Understanding
Obtaining aligned 3D-language data is an inherently difficult task. In 3D object modeling, pioneering works like ShapeNet [14] sourced 3D assets from online repositories, leading to follow-up proliferation of high-quality 3D object datasets [60, 22, 81]. Notably, recent developments include internet-scale data collection with Objaverse [26, 25], accompanied by the integration of object-level captions [83] for 3D-language alignment. Models trained on these datasets demonstrate an enhanced understanding of objects, evident in classification [52], generation [53], and captioning tasks [55].
In contrast, developing datasets for grounded 3D scene understanding is even more challenging due to the extensive requirements for scene acquisition and annotation. Existing works curate RGB-D and scanned indoor scene datasets [23, 13, 78, 9, 67, 58] initially used for benchmarking classical grounding tasks like 3D object detection and segmentation [30, 42, 59, 77, 72]. These semantically labeled scenes are subsequently applied in fine-grained scene grounding tasks like object referral [16, 1, 93], captioning [20, 88, 17, 19], vision-language-navigation [56, 80, 63, 38] and reasoning [5, 57, 37]. Recent work exploits the representation of 3D scene graphs (3DSGs) [4, 79, 69], which concisely describes scenes with hierarchical structures. This representation is notably advantageous for planning [2, 68] and captioning [33], owing to its compatibility with LLMs. Nevertheless, as shown in Tab. 1, these datasets are significantly constrained in both scene and language scales, underscoring the need for scaling up fine-grained scene-language-aligned data to enhance grounded scene understanding.
Vision-Language Learning
Recent years have witnessed tremendous progress in 2D vision-language learning [76, 66, 70, 3, 49, 51, 24], empowered by transformer-based pre-training models [28, 11, 62] and large-scale image-language datasets [15, 71]. A central theme across language and 2D-VL domains is the effectiveness of data scaling [43], as demonstrated by improved alignment and expanded capabilities in open-vocabulary understanding [50, 47, 32, 44] through a simplified contrastive pre-training pipeline [66].
However, in grounded scene understanding, the primary challenge for models has been the limited availability of paired 3D scene-language data, which restricts the application of insights gained from 2D-VL. Current models for 3D scene grounding [35, 94, 86, 6, 54, 40, 41, 18, 82] heavily rely on task-specific knowledge in both model and loss designs or advanced optimization strategies [98]. To bridge this gap, there has been a growing emphasis on employing pre-trained 2D-VL models for 3D-VL [91, 34, 83, 36, 92, 64, 74]. Nonetheless, these models predominantly draw on information available from 2D-VL models (e.g., object attribute, affordance, etc.), falling short on capturing crucial information like object spatial relationships, which are only attainable through 3D data. This urges the need for a multi-level alignment between language and 3D scenes, particularly regarding 3D-specific information. Considering the nascent stage of existing 3D pre-training methods [29, 84, 98], we believe SceneVerse and GPS have the potential to spearhead new avenues in 3D-VL research.
Figure 2: SceneVerse collection and statistics. Given a 3D scene (a), our automated pipeline (c) generates three types of description including scene caption, object caption and object referral. (b) The comparison of different language sources and data composition.
Table 1: Comparison of SceneVerse with existing 3DVL Datasets. SceneVerse expands the data scale of prior work by an order of magnitude. Anno.: human annotations. Syn.: template or LLM generated descriptions.
Dataset 3D Data Language Total
Scene Object Anno. Syn.
ScanRefer[16] |||| |||| 52K - 52K
ReferIt3D[1] 42K 200K 242K
ScanQA[5] 1.5K 33K 27K - 27K
SQA3D[57] |||| |||| - 33K 33K
Multi3DRefer[93] 52K 10K 62K
Cap3D[55] -  [^6] K 58K  [^6] K 724K
ScanScribe[98] 3K 56K 94K 184K 278K
SceneVerse 68K 1.5M 190K 2.3M 2.5M
3 SceneVerse
SceneVerse is the first million-scale dataset designed for grounded scene understanding. Our 3D scenes are curated from diverse existing datasets of both real and synthetic environments. Harnessing the power of 3D scene graphs and LLMs, we introduce an automated pipeline to generate comprehensive and high-quality language for both object-level and scene-level descriptions. We additionally incorporate the most extensive human-annotated object referrals to date, providing new training sources and benchmarks in this field.
3.1 Scene Curation
To address the scarcity of available 3D scene data, we construct SceneVerse by unifying 3D scene data from various existing datasets. We use real-world scene datasets, including ScanNet [23], ARKitScenes [9], HM3D [67], 3RScan [78] and MultiScan [58], alongside synthetic environments from Structured3D [95] and ProcTHOR [27]. The inclusion of these synthetic datasets is mainly motivated by their potential as scalable data sources for 3D-VL alignment. To ensure cohesion across various sources, we conduct preprocessing steps such as room segmentation, point subsampling, axis alignment, normalization, and semantic label alignment. Each scan is represented by a point cloud $\mathrm{P}\in\mathbb{R}^{N\times 8}$, wherein each point is defined by its 3D coordinates, RGB color, instance id and semantic label. In total, we curate $68,406$ 3D scenes in SceneVerse.
3.2 3D Scene Graph Construction
Our 3D scene graph is defined as a set of tuples $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the nodes $\mathcal{V}$ comprises $\mathcal{V}_{1}\bigcup\mathcal{V}_{2}\bigcup\dots\bigcup\mathcal{V}_{K}$, with $\mathcal{V}_{k}$ representing the set of nodes at a particular hierarchical level. Each node $v$ represents one distinct 3D object instance, parameterized by its centroid ${\bm{p}}_{i}\in\mathbb{R}^{3}$ and bounding box size of ${\bm{b}}_{i}=(b_{x},b_{y},b_{z})\in\mathbb{R}^{3}$. The edges $\mathcal{E}$ represent spatial relationships between nodes.
To construct the scene graph $\mathcal{G}$, we first instantiate the nodes with the instance annotation from the point clouds and assign object classes with their corresponding semantic labels. Following prior work[1, 79], we consider the following spatial relations.
Vertical proximity
This encompasses both in-contact relationships (e.g., support, inside, embed), and non-contact ones (e.g., above, below).
Horizontal proximity
Horizontal relationships describe the proximity relations like in front of, next to, behind, etc. Relationships like left, right are contextually dependent on a reference view, where another anchor object is utilized to establish the view direction. The distance between the two objects is also calculated to describe whether the objects are far or near in space.
Multi-object Relationships
This models the spatial arrangement of multiple objects, e.g., align and between.
The node hierarchy is decided by the support relationship. We traverse all the object nodes to calculate spatial relationships, which undergo an automatic verification procedure to rectify incorrect ones. For a more detailed description of the scene graph construction and relationship determination, please refer to Sec. A.2.
3.3 Language Generation with LLMs
The scene-language pairs in SceneVerse aim to capture varying aspects of the 3D scene, which include detailed object attribute descriptions in object captioning, spatial relationships between objects in object referral, and global scene descriptions in scene captioning. Based on the 3D scene graph, we utilize both templates and LLMs to automatically generate descriptions on these three granularities.
Object Captioning
Object captions aim to provide detailed descriptions of an object’s visual and physical properties, facilitating object-level grounding with its distinctive features. Given the multi-view images, we utilize the point cloud of the object $v\in\mathcal{V}$ to identify its occurrence in the images through rendering. The images are then cropped with the rendered bounding boxes and processed through BLIP2 [48] to generate initial object captions. To refine the captions, we select the top 10 sentences with the highest CLIP [66] similarity score and minimal occlusion. The selected sentences are fed into a LLM to obtain a coherent summary of the object captions. In this process, we explicitly instruct the language model to identify and correct the potential errors. The detailed object captioning pipeline is illustrated in Sec. A.3.
Object Referral
[…]
Scene Captioning
The scene-level captions emphasize global information, portraying the key objects in the scene along with their attributes and functionalities. We leverage the constructed 3D scene graph and prompt LLMs to generate these captions. To enhance the diversity of scene captions, we utilize a subgraph sampling strategy, where a subset of edges and nodes are randomly sampled as the scene context. The object counts are also provided as LLM prompts, together with the room type and object attributes if such annotations are available in the dataset.
3.4 Referral Annotation by Humans
In addition to automatically generated scene-text pairs, SceneVerse includes the most comprehensive set of human-annotated, context-rich object referrals to date, serving as a valuable benchmark for assessing grounded scene understanding capabilities. The human annotations contain $96,863$ descriptions in ARKitScenes [9], HM3D [67] and MultiScan [58]. During the annotation process, one human annotator was assigned to write at least 20 words to distinctly refer to a single 3D object within a 3D scene. Each referral text then undergoes independent verification by two additional reviewers, both mandated to accurately locate the referenced object based on the 3D scene and the annotated referral text. Any object referrals that do not pass the verification by either reviewer are flagged for re-annotation.
3.5 Details and Statistics
In total, SceneVerse comprises a total of $68,406$ room-level 3D scans, with the source composition shown in Fig. 2 (b). The dataset contains $1.5$M object instances, comprising $21$ types of relationships following prior work [1, 79]. For the language descriptions, we generate $1$M template-based texts and $1$M sentences by LLM rephrased by Llama [75] and GPT-3.5 [61]. All the rephrasing and summary prompts, along with the complete set of relationships, are detailed in Sec. A.3. To verify the efficacy of our automated language generation pipeline, we conduct a quality check (QC) where 12K generated object-level descriptions are randomly selected for human verification, achieving a $96.93\%$ pass rate. This shows the capability of our proposed scene-graph-based generation approach to produce high-quality language descriptions, laying a robust foundation for future scalability.
4 Grounded Pre-training for Scenes
Figure 3: Overview of our proposed GPS model. We leverage contrastive alignment in three levels $\mathcal{L}_{\text{obj}}$, $\mathcal{L}_{\text{scene}}$, and $\mathcal{L}_{\text{ref}}$ as well as a masked language modeling objective $\mathcal{L}_{\text{MLM}}$ for model learning.
In this section, we introduce GPS, an efficient transformer-based model trained with multi-level contrastive losses for aligning 3D scenes and texts. As shown in Fig. 3, we echo the language descriptions collected at different levels to form scene-language pairs at both object-level, referral-object-level, and scene-level for contrastive objectives in GPS. We describe the design of each level in the following sections.
4.1 Object-level Grounding
Given a 3D scene point cloud $\mathcal{S}$, we use an off-the-shelf 3D object segmentation model to decompose it into a bag of $N$ objects $\mathcal{S}=\left\{{\bm{o}}_{1},{\bm{o}}_{2},\cdots,{\bm{o}}_{n}\right\}_{i=1}^{N}$. We extract object features $\{{\bm{f}}^{O}_{i}\}$ with an object point cloud encoder and text features $\{{\bm{f}}^{T}_{i}\}$ by feeding object-captions $\{{\bm{T}}^{\text{obj}}_{i}\}$ into a frozen language model. Following [83], we perform cross-modal alignment on the object features and text features via:
$\mathcal{L}_{\text{obj}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(p,r)\right)}}+\right.$ (1)
$\left.\ \ \ \log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(r,q)\right)}}\right),$
where $D^{\text{obj}}(p,q)=({\bm{f}}^{O}_{p}{\bm{f}}^{T}_{q}/\tau)$ denotes the dot product between object and text features and $(p,q)$ denotes a pair of aligned object-text pair in the training batch and $r$ iterates over all object-text pairs in the training batch. Similar to CLIP [66], we use a learnable temperature parameter $\tau$ to facilitate model learning.
4.2 Scene-level Grounding
With aligned object features, we encode the scene by incorporating object spatial locations into the extracted object features. Specifically, we use a spatial transformer model to encode extracted object features $\{{\bm{f}}^{O}_{i}\}$ with their spatial location features $\{{\bm{l}}_{i}\}$ following [18, 98]:
${\bm{f}}^{S}=\mathrm{SpatialAttn}(\{{\bm{f}}_{i}^{O}\},\{{\bm{l}}_{i}\})$
where $\{{\bm{f}}_{i}^{S}\}$ denotes the feature of object ${\bm{o}}_{i}$ after encoding with spatial location features. To perform scene-level alignment, we operate on these scene-level object features $\{{\bm{f}}_{i}^{S}\}$ and align it with the scene caption ${\bm{T}}^{\text{scene}}$. Specifically, we feed the object features into a projection layer and use max-pooling over all object features to obtain the scene feature ${\bm{g}}^{S}$. Similar to object-level grounding, we pass the scene caption through a tunable language model to obtain text feature ${\bm{g}}^{T}$ and perform scene-level contrastive alignment through:
$\mathcal{L}_{\text{scene}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(p,r)\right)}}+\right.$ (2)
$\left.\ \ \ \log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(r,q)\right)}}\right),$
where $D^{\text{scene}}(p,q)=({\bm{g}}_{p}^{S}{\bm{g}}_{q}^{T}/\tau)$ denotes the dot product between scene feature ${\bm{g}}_{p}^{S}$ and scene caption feature ${\bm{g}}_{q}^{T}$ for each pair of aligned scene-text pairs in the training batch and $r$ iterates over all scene-text pairs in the training batch.
4.3 Referral-object-level Grounding
To model the relationships revealed in referring expressions, we employ a self-attention-based reasoning transformer for grounding object referrals in scenes. This transformer takes in scene-object features $\{{\bm{f}}^{S}_{i}\}$ and an object referral ${\bm{T}}^{\text{ref}}$ and performs self-attention to learn relationships between text descriptions and object relationships. We use the same tunable language encoder as in scene-level grounding for extracting per-object referral features. We pass this text feature together with scene-object features into the self-attention transformer to obtain the aligned object features ${\bm{h}}^{S}_{i}$ and the sentence-level referral feature ${\bm{h}}^{T}$. We then perform the referral-object-level contrastive alignment following:
$\mathcal{L}_{\text{ref}}=-\log\frac{\exp{\left(\bar{{\bm{h}}}^{S}{\bm{h}}^{T}/\tau\right)}}{\sum_{p}\exp{\left({\bm{h}}_{p}^{S}{\bm{h}}^{T}/\tau\right)}},$ (3)
where $\bar{{\bm{h}}}^{S}$ denotes the feature of the referred object, $p$ iterates over all objects within the same scene. Notably, in contrast to inter-scene contrast that was done in object- and scene-level alignment, we force the selection of positive pairs to be within the same scene to provide intra-scene contrast for fine-grained object grounding. This mimics the success of intra-image and inter-image contrasts commonly used for region-word alignment in 2D-VL models [90].
To learn the multi-level alignment between 3D scenes and language, we first train the point cloud encoder with object-level grounding objects to obtain a good feature initialization for grounding objects in scenes. During the scene grounding stage, we train our inter- and intra-scene objectives together with a mask language modeling loss $\mathcal{L}_{\text{MLM}}$ over the inputted object-referral texts to tune the parameters within the language encoder and self-attention transformer. Above all, the learning of GPS could be summarized as optimizing the following objective:
$\mathcal{L}=\mathcal{L}_{\text{obj}}+\mathcal{L}_{\text{scene}}+\mathcal{L}_{\text{ref}}+\mathcal{L}_{\text{MLM}}.$
5 Experiments
In this section, we present the evaluation results addressing the following questions:
  • How effective is the data scaling in SceneVerse for 3D visual grounding? Does the scale-up work for general pre-training based 3D-VL models?
  • How well is the GPS pre-training pipeline? Does it exhibit similar properties of 2D-VL models in 3D-VL tasks?
  • What potentials are offered by SceneVerse and GPS for future research? What is missing?
In the following sections, we describe in detail about model performance regarding these key topics. Due to the page limit, we direct readers to the Appendices B and C for implementation details and more experimental analyses.
5.1 3D Visual Grounding
Table 2: 3D Visual Grounding results on Nr3D, Sr3D, and ScanRefer. We use “direct” for our model trained on SceneVerse with no additional fine-tune head, and “fine-tune” for the data-specific fine-tuned version of our model. We highlight the best results in bold.
Method Nr3D Sr3D ScanRefer Acc@0.5
Overall Easy Hard V-Dep. V-Indep. Overall Easy Hard V-Dep. V-Indep. Overall Unique Multiple
3DVG-Trans [94] 40.8 48.5 34.8 34.8 43.7 51.4 54.2 44.9 44.6 51.7 34.7 60.6 28.4
TGNN [39] 37.3 44.2 30.6 35.8 38.0 45.0 48.5 36.9 45.8 45.0 29.7 56.8 23.2
TransRefer3D [35] 48.0 56.7 39.6 42.5 50.7 57.4 60.5 50.2 49.9 57.7 - - -
InstanceRefer [89] 38.8 46.0 31.8 34.5 41.9 48.0 51.1 40.5 45.8 48.1 32.9 66.8 24.7
FFL-3DOG [31] 41.7 48.2 35.0 37.1 44.7 - - - - - 34.0 67.9 25.7
LAR [6] 48.9 58.4 42.3 47.4 52.1 59.4 63.0 51.2 50.0 59.1 - - -
SAT [86] 56.5 64.9 48.4 54.4 57.6 57.9 61.2 50.0 49.2 58.3 30.1 50.8 25.2
3D-SPS [54] 51.5 58.1 45.1 48.0 53.2 62.6 56.2 65.4 49.2 63.2 37.0 66.7 29.8
3DJCG [12] - - - - - - - - - - 37.3 64.3 30.8
BUTD-DETR [41] 54.6 60.7 48.4 46.0 58.0 67.0 68.6 63.2 53.0 67.6 39.8 66.3 35.1
MVT [40] 59.5 67.4 52.7 59.1 60.3 64.5 66.9 58.8 58.4 64.7 33.3 66.5 25.3
ViL3DRel [18] 64.4 70.2 57.4 62.0 64.5 72.8 74.9 67.9 63.8 73.2 37.7 68.6 30.7
EDA [82] 52.1 58.2 46.1 50.2 53.1 68.1 70.3 62.9 54.1 68.7 42.3 68.6 37.6
3D-VisTA (scratch[98] 57.5 65.9 49.4 53.7 59.4 69.6 72.1 63.6 57.9 70.1 41.5 70.9 34.8
3D-VisTA [98] 64.2 72.1 56.7 61.5 65.1 76.4 78.8 71.3 58.9 77.3 45.8 75.1 39.1
Ours (scratch) 58.7 67.0 50.9 55.8 59.8 68.4 70.5 63.4 53.1 69.0 40.4 71.3 34.7
Ours (pre-train) 55.2 62.8 48.0 45.5 58.8 74.1 76.4 68.5 54.1 75.0 47.1 77.4 41.6
Ours (fine-tuned) 64.9 72.5 57.8 56.9 67.9 77.5 80.1 71.6 62.8 78.2 48.1 77.9 42.7
Settings
We evaluate our model on three commonly-used datasets for 3D visual grounding: ScanRefer [16], Nr3D, and Sr3D [1]. For Nr3D and Sr3D, we follow Achlioptas et al. [1] and report the grounding accuracies of models using ground-truth object masks. For ScanRefer, we follow Zhu et al. [98] and use Mask3D [72] to generate object proposals. Results are reported as $\text{Acc}@0.5$ to evaluate the correctness of predictions whose object bounding boxes overlap the ground truth with IoU $>0.5$. For comparisons, we compare with existing baselines by providing the results of pre-trained GPS and dataset-specific fine-tuned GPS. Please see more details in the Appendix C.
Results and Analyses
As shown in Tab. 2, GPS trained on SceneVerse achieves state-of-the-art results on all existing 3D-VL grounding benchmarks. Initially, when GPS is trained directly on the training sets of benchmark datasets, labeled as Ours (scratch), it underperforms compared to existing models that employ more complex structures or loss designs. This result underscores the data-intensive nature of the contrastive alignment paradigm. However, when presented with extensive training data in SceneVerse, the results of our model without additional fine-tuning, i.e., Ours (pre-train), significantly improves and already achieves state-of-the-art results on benchmarks like ScanRefer. Moreover, the dataset-specific fine-tuned model, i.e., Ours (fine-tuned), consistently outperforms existing baselines with only a simple projection MLP added on top of the pre-trained model, jointly optimized during fine-tuning without any other auxiliary architecture or loss objective. These results underscore the strong potential of both the SceneVerse and GPS for 3D-VL tasks.
5.2 Zero-Shot Transfer
Settings
To better evaluate the effectiveness of both the SceneVerse data and the GPS model, we further perform zero-shot transfer experiments to test the models’ capability in 4 benchmarks, ScanRefer, Sr3D, Nr3D, and SceneVerse-val. We create SceneVerse-val using $8.5K$ annotated object referrals of $271$ scenes in MultiScan, and randomly split the scenes following a 4:1 train / test split for creating the held-out test set. We mainly consider 2 specific transfer settings in our experiments: (i) zero-shot: models trained by removing all the scenes from the target dataset, tested on held-out unseen scenes, and (ii) zero-shot text: Models trained on data that include the 3D scenes from training set of the target dataset, yet tested exclusively with unseen scene-text distribution. Specifically, for the zero-shot text setting, we use the generated texts in SceneVerse as fine-tuning sources for the zero-shot model. We mainly compare our model against a recent pre-training-based model 3D-VisTA. See more details on experimental setting and implementation in the Appendix C.
Table 3: Zero-shot transfer results on established benchmarks.
Method Nr3D Sr3D ScanRefer@0.25 ScanRefer@0.5
3D-VisTA (scratch) 57.5 69.6 45.9 41.5
3D-VisTA (zero-shot) 35.2 31.2 33.2 29.6
3D-VisTA (zero-shot text) 43.1 36.1 41.1 36.4
Ours (scratch) 58.7 68.4 44.5 40.4
Ours (zero-shot) 32.4 33.3 35.2 31.1
Ours (zero-shot text) 41.9 38.1 40.7 35.8
Table 4: Zero-shot transfer on SceneVerse-val. We evaluate models following settings in Nr3D/Sr3D using GT object proposals.
Method Overall Easy Hard V-Dep. V-Indep.
3D-VisTA (scratch) 40.7 53.1 21.6 37.3 44.3
3D-VisTA (zero-shot) 52.9 59.6 35.4 53.7 52.2
3D-VisTA (zero-shot text) 58.1 70.0 39.6 52.5 64.1
Ours (scratch) 38.5 50.2 20.8 33.7 43.9
Ours (zero-shot) 59.2 69.4 44.0 53.1 66.3
Ours (zero-shot text) 60.6 70.9 45.1 54.8 67.3
Results and Analyses
We present the results of zero-shot transfer experiments in Tab. 3 and Tab. 4 with the following key observations:
  • Our GPS model demonstrates superior generalization to unseen scenes compared to the 3D-VisTA model. In zero-shot transfer scenarios, our model consistently outperforms 3D-VisTA across established benchmarks and SceneVerse-val. This indicates the effectiveness of contrastive alignment over traditional classification objectives, aligning with the advancements seen in 2D-VL models for open-vocabulary grounding and transfer capabilities
  • SceneVerse dataset substantially enhances 3D-VL grounding capabilities through zero-shot transfer, especially when provided with relatively limited training data, i.e., SceneVerse-val. As demonstrated in  Tab. 4, there is a significantly improved performance when comparing models trained on SceneVerse in a zero-shot manner to those trained from scratch. This indicates that SceneVerse can effectively capture knowledge for general 3D scene grounding. Consequently, this underscores its potential as a go-to pre-training dataset for 3D-VL tasks.
  • The impact of our extensive collection and scalable generation of scene-text pairs is further evidenced by the results in the zero-shot text setting. Notably, as shown in Tab. 3, these automatically generated scene-text pairs supply ample knowledge for comprehending the scene distribution. This contributes significantly to the substantial improvement over the zero-shot performance.
Figure 4: Model performance v.s. data scale. Models consistently improve in both the pre-train and zero-shot transfer settings on ScanRefer and SceneVerse-val with data scaling-up.
5.3 Ablative Studies and Discussions
In this section, we present ablative studies focused primarily on the data collected in SceneVerse. Our aim is to elucidate the effects of data scaling more clearly. For discussions regarding ablation studies on model architecture, readers are referred to the Appendix D. The following points are specifically discussed in this section.
How important is data-scaling?
We conduct ablation studies over the amount of data used while pre-training GPS. We consider the model trained with $\frac{1}{8}$, $\frac{1}{4}$, $\frac{1}{2}$ of SceneVerse to show the effectiveness of data-scaling on model performance in the pre-train and zero-shot transfer settings in ScanRefer and SceneVerse-val. As shown in Fig. 4, we observe consistent performance improvement over the increase of data scale for both settings. We provide additional experiments in the Appendix D to show that such scaling effect is not only beneficial for 3D-VL grounding but also for other 3D tasks like semantic segmentation [72, 85].
How is the generated data compared with human-annotated data?
We assess the performance of models trained using various scene-text sources, specifically focusing on their performance in the ScanRefer dataset without additional fine-tuning. As shown in Tab. 5, models trained with our template-based generated texts and Large Language Model (LLM)-refined texts show significant improvements over models trained solely on ScanRefer. More importantly, these variants of our model already achieve state-of-the-art results compared with previous baselines. This indicates the effectiveness of our text-generation pipeline. Finally, we observe that adding human-annotated data is still beneficial for model performance. However, the improvement is relatively marginal over models trained on our generated data.
What is the role of the synthetic scenes in this scale-up process?
With synthetic data providing large-scale and diverse scene data for 3D-VL tasks, we evaluate the models’ domain transfer (Sim2Real) capability. Specifically, we compare models trained on all real scenes in SceneVerse against models trained exclusively on two synthetic subsets of SceneVerse, i.e., Structured3D and ProcTHOR. As shown in Tab. 6, models trained on synthetic subsets demonstrate remarkable performance on their corresponding test sets while suffering when transferred to real or other synthetic scenes. In contrast, the model trained on real scene-text pairs exhibits less severe performance drops when generalizing to synthetic scenes. This result affirms the domain gap between real and synthetic scenes in 3D-VL grounding and shows that a simple scale-up in the amount of scenes is insufficient when the scene naturalness can not be guaranteed. Considering the scalability of our quality-ensured language generation and also the scaling effect shown in our experiments, the rate-determining step for further scaling-up 3D-VL comes to the collection of diverse, high-quality, and realistic scenes that capture natural 3D scene distributions.
Table 5: Ablation on scene-text pair types used in training. We report model results on ScanRefer with no additional finetuning.
Template LLM Anno. Acc@0.25 Acc@0.5
43.5 38.4
50.9 46.1
51.1 46.3
52.0 47.1
Table 6: Cross domain transfer results of models learned in real and synthetic datasets without additional fine-tuning. “S3D” stands for Structured3D.
Real Synthetic SceneVerse-val S3D ProcTHOR
All 64.8 37.1 43.4
S3D 7.0 85.1 16.1
ProcTHOR 4.2 16.3 91.0
6 Conclusion
In this work, we scale up 3D-VL in the context of grounded scene understanding. We introduce SceneVerse, a million-scale 3D-VL dataset encompassing various scenes and multi-level scene descriptions sourced from both human annotation and our proposed scene-text generation approach. Utilizing SceneVerse, we propose Grounded Pre-training for Scenes, a model trained with multi-level scene-language contrastive alignment over the data collected. Through extensive experiments, we show that GPS achieves state-of-the-art results on all existing 3D-VL grounding tasks. We further conduct zero-shot transfer experiments to show the improved generalization performances of GPS trained on SceneVerse compared with previous baselines. We hope our efforts and successful scale-up attempts in SceneVerse could pave the way for a new research paradigm in 3D-VL.
7 Acknowledgement
The authors thank Yaowei Zhang from BIGAI for designing the result visualization framework, Jiangyong Huang and Xiongkun Linghu from BIGAI for suggestions on data generation and refinement, and colleagues from BIGAI for their helpful discussions and suggestions.
Supplementary Material
In Appendix A, we introduce more details of SceneVerse, including the 3D scene preprocessing, scene graph construction and automatic language generation. Appendix B presents more model and implementation details. Appendix C include a more in-depth summary of the settings and implementations for the experiments in the main paper, as well as the ablative study and additional experiments on semantic segmentation to demonstrate the benefits of SceneVerse.
Appendix A SceneVerse
A.1 3D Scenes
To address the scarcity of available 3D scene data, we construct SceneVerse by unifying 3D scene data from various existing datasets. The curation involves utilizing real-world scene datasets such as ScanNet [23], ARKitScenes [9], HM3D [67], 3RScan [78] and MultiScan [58],in conjunction with synthetic environments from Structured3D [95] and ProcTHOR [27]. The incorporation of these synthetic datasets is primarily driven by their potential as scalable data sources for 3D-VL alignment. To facilitate the training process, we conduct the following preprocessing steps.
Room Segmentation
The 3D scenes in HM3D and ProcTHOR are released at the building level, encompassing multiple rooms and sometimes spanning over 50 meters. To align with existing benchmarks [16, 1], we leverage the associated metadata to segment the 3D point cloud at the room level, facilitating subsequent operations in scene graph construction and language description generation. Additionally, we implement a filtering process to exclude extremely large rooms and those with fewer than 4 objects in the scene.
Point Cloud Normalization
In order to mitigate the data disparities arising from diverse capture devices across various data sources, we subsample each point cloud to a maximum of $240,000$ points. Each point cloud then undergoes a transformation centered on the central point on the floor, followed by rotation to align the room layout with the axis following the approach by Chen et al. [18].
Semantic Label Alignment
Given the divergence in semantic label sets across different datasets, we undertake a comprehensive effort to map all the object class labels to the $607$ semantic labels in ScanNet [23] to facilitate close-vocabulary object classification [65] in the existing model framework [98]. We construct the mapping in each dataset through LLM and manual verification. Note that the object-level grounding in GPS can directly deal with open-set object labels or captions, similar to CLIP [36].
After the preprocessing, each scan is represented by a point cloud $\mathrm{P}\in\mathbb{R}^{N\times 8}$, wherein each point is defined by its 3D coordinates, RGB color, instance id and semantic label. In total, we curate $68,406$ 3D scenes in SceneVerse.
A.2 3D Scene Graph Construction
In Sec. 3.2, we introduce an automated pipeline to construct 3D scene graphs from point clouds. Here, we provide more implementation details and the relationship definition.
A.2.1 Relationships
Table A.1: Relationships in SceneVerse. The 3D scene graph captures 21 types of relationships ranging in 4 categories.
Category Relation
|||| supported by
In-contact vertical embedded into
placed in
|||| inside
hanging on
affixed on
mounted on
Non-contact vertical above
higher than
below
lower than
near(far) to the left of
near(far) to the right of
is behind
Horizontal is in front of
close to
adjacent to
besides
next to
Multi-object between
aligned
Our 3D scene graph captures $21$ types of relations as shown in Tab. A.1. We provide illustrations of how these relations are defined in the 3D space, as can be seen in Fig. A.1.
A.2.2 Scene Graph Construction
Due to the inherent noise and incompleteness in the point cloud representation, automatically extracting precise and comprehensive relationships from the point clouds is a non-trivial task. Below we detail our 3D scene graph construction process, as outlined in Alg. 1.
We first instantiate the graph nodes with the instance annotation from the point cloud and parameterize each node with object centroid $\mathrm{p_{i}}\in\mathbb{R}^{3}$ and size of the axis-aligned bounding box $\mathrm{b_{i}}=(b_{x},b_{y},b_{z})\in\mathbb{R}^{3}$ (Line 1-3). Next, we traverse all the nodes to determine their spatial relationships (Line 4-22). Notably, in cases where an object node lacks any in-contact vertical relationships with other objects in the scene, we designate such objects as "hangable" and calculate their non-contact vertical relationships (Line 9-13). Examples of such objects include paintings, curtains, etc. Finally, we establish relationships between multiple objects (Line 23): i) When a target object is connected with two edges labeled left and right, the target object, along with the two neighboring nodes, forms a between relationship triplets. ii) If the offset of the center point coordinates of a group of objects in either the X-axis or Y-axis direction is smaller than a specified offset threshold $\delta$, then this group of objects forms an align relationship. The offset threshold $\delta$ will be adjusted based on the size of the scene. In additional, we utilize an automatic verification procedure to validate the scene graph, further improving the quality of the scene graph we constructed (line 24). One of the verification operations involves manually maintaining a mapping between objects and relationship descriptions based on common sense. For example, people usually use “mounted on” to describe the relation between TV and wall, rather than “hanging on”. Therefore, we would automatically refined ( TV, hanging on,  wall) to ( TV, mounted on,  wall).
In our constructed 3D scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the nodes $\mathcal{V}$ comprises $\mathcal{V}_{1}\bigcup\mathcal{V}_{2}\bigcup\dots\bigcup\mathcal{V}_{K}$, with $\mathcal{V}_{k}$ representing the set of nodes at a particular hierarchical level. The hierarchies are determined by the support relationship; for instance, objects supported by the floor constitute $\mathcal{V}_{0}$, while objects supported by the table will form $\mathcal{V}_{1}$, etc. Note that edges originating from one node $v\in\mathcal{V}_{k}$ may only terminate in nearby hierarchies $\mathcal{V}_{k}\cup\mathcal{V}_{k+1}\cup\mathcal{V}_{k+1}$. In other words, edges in the scene graph exclusively connect nodes within the same hierarchical level, or one level higher or lower.
Input :  $M$ object point clouds $\{P_{1},P_{2},\dots,P_{m}\}$
Output : 3D scene graph $\mathcal{G}(\mathcal{V},\mathcal{E})$
1:   for $i$ from $1$ to $M$ do
2:      Create node $v_{i}\in\mathcal{V}$ using the centroid $p_{i}$ and bounding box size $b_{i}$ of object point cloud $P_{i}$
3:   end for
4:   for $i$ from $1$ to $M$ do
5:      for $j$ from $i+1$ to $M$ do
6:         $\mathrm{RelsType}_{v}\leftarrow\mathrm{VerticalInContact}(v_{i},v_{j})$
7:         Add in-contact vertical relationship triplets $(v_{i},v_{j},e_{i,j})$ with $\mathrm{RelsType}_{v}$ to $\mathcal{G}$
8:      end for
9:      if No objects horizontally related to $v_{i}$ then
10:         for $k$ from $1$ to $M$ and $i\neq k$ do
11:            $\mathrm{RelsType}_{v}\leftarrow\mathrm{VerticalNonContact}(v_{i},v_{k})$
12:            Add non-contact vertical relationship triplets $(v_{i},v_{k},e_{i,k})$ with $\mathrm{RelsType}_{v}$ to $\mathcal{G}$
13:         end for
14:      end if
15:   end for
16:   for  $v_{i}\in\mathcal{V}$ do
17:      let ${\{v_{i_{1}},v_{i_{2}},...,v_{i_{N}}\}}$ be the $N$ different nodes with the same in-contact vertical parent node $v_{i}$
18:      for $j$ from $1$ to $N$ do
19:         $\mathrm{RelsType}_{h}\leftarrow\mathrm{Horizontal}(v_{i},v_{i_{j}})$
20:         Add horizontal relationship triplets $(v_{i},v_{i_{j}},e_{i,i_{j}})$ with $\mathrm{RelsType}_{h}$ to $\mathcal{G}$
21:      end for
22:   end for
23:   Update $\mathcal{G}\leftarrow\mathrm{MultiObjects}(\mathcal{G})$
24:   Update $\mathcal{G}$ with automatic verification procedure
Algorithm 1 Scene Graph Construction Pipeline
Figure A.1: Overview of the relationships in SceneVerse. The target object is colored in blue.
Figure A.2: Examples of object captioning. We color the target object in bold.
Figure A.3: Examples of scene captioning.
Figure A.4: Examples of object referral. Note that the green bounding box indicates the target object and yellow bounding box indicates the anchor object(s).
A.3 Language Generation Details
In Sec. 3.3, we adopt both templates and LLM to automatically generate scene-language pairs in SceneVerse. More technical details and examples are provided in this section.
A.3.1 Object Captioning Pipeline
Object captions aim to provide detailed descriptions of an object’s visual and physical properties, facilitating object-level grounding with its distinctive features. The detailed object captioning pipeline is outlined in Alg. 2. Given the multi-view images $\{I_{1},I_{2},\dots,I_{n}\}$, we utilize the point cloud $P_{o}$ of the object $o$ to get the visible points $P_{o,v}^{vis}$ in the images $v$ through rendering. The occlusion score $s_{o,v}^{occ}$ is calculated as the ratio between the number of visible points and the object point cloud. The image is then cropped with the rendered bounding box and processed through BLIP2 [48] to generate the initial object caption $C_{o,v}$. For each initial caption, we calculate its CLIP [66] similarity score between the text and the cropped image, denoted by $s_{o,v}^{clip}$. To get a refined object caption, we select the top $10$ initial captions with the highest CLIP score and minimal occlusion. The selected sentences are fed into a LLM to obtain a coherent summary of the object captions. In this process, we explicitly instruct the language model to identify and correct the potential errors.
Input :  $M$ object point clouds $\{P_{1},P_{2},\dots,P_{m}\}$; $N$ multiview images $\{I_{1},I_{2},\dots,I_{n}\}$
Output : Captions for each object in the scene $\{C_{1},C_{2},\dots,C_{m}\}$
1:   for $o=1,2,\dots,M$ do
2:      for $v=1,2,\ldots,N$ do
3:         Project $P_{o}$ on $I_{v}$ to get visible points $P^{vis}_{o,v}$
4:         Crop $I_{v}$ with the bounding box of $P^{vis}_{o,v}$ to get $I^{crop}_{o,v}$
5:         Get the image caption $C_{o,v}$ for $I^{crop}_{o,v}$ using BLIP2 [48]
6:         Calculate the similarity score $s^{clip}_{o,v}$ between $C_{o,v}$ and $I^{crop}_{o,v}$ with CLIP [66]
7:         Calculate the occlusion score $s^{occ}_{o,v}=\frac{\#P^{vis}_{o,v}}{\#P_{o}}$
8:      end for
9:      Select the top-10 $\{C_{o,v}\}$ with highest $s^{clip}_{o,v}*s^{occ}_{o,v}$
10:      Summary selected $\{C_{o,v}\}$ with GPT-3.5 to get $C_{o}$
11:   end for
Algorithm 2 Object Captioning Pipeline
A.3.2 Automatic Language Generation
Template-based
We create diverse templates to generate descriptions for each type of relationship. We categorized the templates into three types based on the number of objects involved and their spatial relationships.
  • Pair-wise: The pair-wise templates are used to describe the positional relationship between the target object and the anchor object in the scene. We design various templates to enrich the templated-based descriptions, spanning active and passive tense, as well as inversion clauses. Typical examples are shown below:
      The target-object (is) spatial-relation the anchor-object.
      It is a target-object that (is) spatial-relation the anchor-object.
      There is a target-object that (is) spatial-relation the anchor-object.
      Spatial-relation the anchor-object is the target-object.
      Spatial-relation the anchor-object, a target-object is placed.
  • Multi-objects: This is utilized when the target object forms a between or align relationship with multiple anchor objects in the scene. The templates follow the same construction rules as the Pair-wise templates.
  • Star-reference: To increase complexity in templated-based descriptions, we design “star-reference” to describe the target object and its relationship with 3 randomly selected anchor objects in the scene graph. In particular, we perform cluster analysis on the selected relationship triplets. Based on the diversity of the analysis, different templates will be chosen to generate descriptions. For example, when the relations between 3 anchor objects and the target object is the same, we prefer to use the template like: “The target-object (is) spatial-relation the anchor-object-1anchor-object-2 and anchor-object-3”. If 2 out of the 3 anchor objects have the same relations with the target object, we would use a template like: “The target-object (is) spatial-relation-1 the anchor-object-1 and anchor-object-2, and (is) spatial-relation-2 the anchor-object-3.”
Table A.2: Prompts used in SceneVerse.
Description type
Prompt
Object caption
Summarize caption below. The summary should be a description of the target-object. Focus on the target-object’s attribute, like color, shape and material, etc. Identify and correct the potential errors.
caption: A bed in a hotel room. A white comforter on a bed. A bed with a striped comforter…
target-object: Bed
Object referral
Rewrite the following caption using one random sentence structure. You should give me only one rewritten sentence without explanation.
caption: The bed is between desk and nightstand.
Rewrite the following caption. You should give me only one rewritten sentence about target-object without explanation. Make sure target-object is the subject of the sentence, not anchor-object(s). If the sentence is in full inversion, keep the inversion.
caption: The armchair is next to the sofa.
target-object: Armchair
anchor-object(s): Sofa
Rewrite the following caption using one random sentence structure. You need to focus on the location and relations of the target-object that appears in the sentence. If multiple target-object appear in the sentence, you need to focus on the first target-object that appears. You can also add the target-object’s function and comfort level based on the sentence, e.g., how the objects can be used by humans and human activities in the scene. You should give me only one rewritten sentence without explanation.
caption: Far from the bowl and peppershaker, the vase is to the left, it is also on the top of countertop.
target-object: Vase
Scene captioning
Your task is to provide a summary for a scene from a given scene graph. The scene contains some objects, which compose a scene graph in json format.
There are 3 types of descriptions in scene graph: “scene type” denotes the type of the scene. “objects count” then listed the objects in the scene and their quantity, it should be noted that the actual objects in the room may be more than listed. “objects relations” describe the spatial relations with objects.
Also describe the scene concerning commonsense, e.g., how the objects can be used by human and human activity in the scene. The description should conform to the given scene graph. The spatial relations between objects can only be inferred from the “objects relations“ in scene graph. Don’t describe each object in the scene, pick some objects of the scene for summary. Don’t describe each relations in the scene, pick some relations of the scene for summary. You can also summarize the room’s function, style, and comfort level based on the arrangement and count of objects within the room. The summary should be about the object types, object attributes, relative positions between objects. Your summary must not exceed 80 words. You must write using one random sentence structure.
scene graph: { ’scene_type’: ’Bedroom’, ’object_count’: {’nightstand’:2, ...}, ’relation’: {’nightstand’, ’on’, ’floor’}, {’backback’, ’in front of’, bed}, ...}
LLM-rephrasing
To increase description diversity we use the GPT-3.5[61] and Llama[75] for description rephrasing. This improves the diversity and naturalness of the template-based descriptions, as is shown in Fig. 2. The detailed prompts are provided in Tab. A.2.
More examples of the scene-language pairs in SceneVerse are shown in Fig. A.2, Fig. A.3 and Fig. A.4.
Appendix B Model Details
B.1 Spatial-Attention Transformer
In Sec. 4.2, we leveraged and spatial-attention based transformer architecture to aggregate object-level point cloud features with spatial location information. In this section, we provide the detailed design of this proposed module.
Formally, given object features $\{{\bm{f}}_{i}^{O}\}_{i=1}^{N}$ and their locations $\{{\bm{l}}_{i}\}_{i=1}^{N}$, we first construct pair-wise spatial relationship feature via:
${\bm{m}}_{ij}=\left[d_{ij},\sin(\theta_{h}),\cos(\theta_{h}),\sin(\theta_{v}),\cos(\theta_{v})\right],$
where $d_{ij}$ denotes the Euclidean distance between objects and $\theta_{h},\theta_{v}$ are the horizontal and vertical angles of the line connecting the centers of objects $i$, $j$. We then use this pair-wise relationship feature $M=[{\bm{m}}_{ij}]\in\mathbb{R}^{N\times N\times 5}$ to modulate the attention weights of the self-attention layer in the transformer when aggregating object features as shown below:
$\small\mathrm{Attn}(Q,K,V,M)=\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{d_{h}}}+\log\sigma(M\omega)\right)V,$
where $\omega\in\mathbb{R}^{5}$ is a projection layer mapping spatial pair-wise features to the attention scores and $\sigma$ denotes the sigmoid function. This process could be equivalently interpreted as using the spatial location of objects to adjust the self-attention feature aggregation between objects, making spatially related objects have more attention weights.
B.2 Pre-training Details
For training our model GPS, we conduct a two-stage training approach. As described in Sec. 4.3, we first pre-train the object point cloud encoder with the object-level grounding objective. Next, we freeze the object point cloud encoder during the second pre-training stage for scene-level pre-training that includes model training with scene-level grounding and referral object grounding objectives. This design is inspired by recent works like [98, 21] that demonstrated a well-grounded initialization of object representations is beneficial for 3D scene grounding.
Object-level pre-training
To correctly align objects in scenes with their captions, we utilize the ground-truth object bounding boxes provided with the datasets to segment all objects in the scenes. Next, we utilize a PointNet++ [65] encoder to encode and align these object point clouds with object captions provided in SceneVerse following Sec. 4.1. For object instances with no object captions synthesized, we follow [66] and construct captions with their semantic class labels like “the point cloud of <CLASS>”. Notably, as our model design sets no constraints on object point cloud encoders, the choice of object encoder mainly depends on the computing resources available.
Scene-level pre-training
With pre-trained object feature extractors, we further use both scene captions and object-referring expressions for scene-level pre-training. We use a 4-layer BERT encoder for encoding both scene captions and object referrals. As discussed in Sec. 4.2, we apply a 4-layer spatial transformer to encode object features with their locations. For scene-level grounding, we adopt a max-pooling layer to aggregate features from the spatial transformer and align with the [CLS] feature of the scene caption. For referral-object-level grounding, we further pass the obtained object features as well as the referral language features into a 4-layer self-attention transformer and use the grounding objective described in Sec. 4.3 to match the referred object’s feature and the [CLS] feature of the referring expression.
Training
For object-level pre-training, we utilize an AdamW optimizer with a learning rate of $1\times 10^{-2}$ for 1500 epochs and no warm-up periods. During training, we use a batch size of 512 and leverage a cosine annealing scheme for learning rate scheduling with a minimum learning rate of $1\times 10^{-3}$. For scene-level pre-training, we use an AdamW optimizer with a learning rate of $1\times 10^{-5}$ for the language encoder, a learning rate of $1\times 10^{-4}$ for the spatial transformer, a learning rate of $1\times 10^{-4}$ for the self-attention transformer, and a learning rate of $5\times 10^{-4}$ for all remaining learnable parameters (e.g., projections). For all experiments, we train the model for 150 epochs with a warm-up period of 500 and also a cosine annealing scheme for learning rate with a minimum learning rate ratio of $0.1$. All pre-training experiments are run on 8 NVIDIA-A100 GPUs with the longest pre-training on SceneVerse taking about 2 days.
Appendix C Experimental Details
In this section, we provide details on experimental settings, model implementation, and additional results.
C.1 3D Visual Grounding
Setting
For all datasets, we evaluate all models with only the training sets provided. Following previous works [98], we report model performance on the validation set of all datasets in Tab. 2. Notably, we used an off-the-shelf Mask3D segmentation model for generating object proposals with no optimization.
Implementation
As briefly mentioned in Sec. 5.1, we mainly considered three model settings in 3D visual grounding experiments, namely scratch, pre-train, and fine-tuned. For the pre-train setting, we follow the same setting mentioned in Sec. B.2. In the scratch and fine-tuned settings, to fairly compare with other dataset-specific fine-tuned models, we add an additional 2-layer MLP over the object features from the referral grounding self-attention transformer. During training, we fine-tune this grounding head together with all model weights for 100 epochs with a learning rate of $1\times 10^{-4}$ for the added projection layer and set all other settings the same as the implementation described in Sec. B.2.
C.2 Zero-shot Transfer
Setting
In the zero-shot experiments, we first construct the held-out test set by aggregating scene-text pairs in SceneVerse from scenes in ScanNet and MultiScan. Specifically, we use the validation set of ScanRefer, Nr3D, and Sr3D. For scene-text pairs in the SceneVerse-val, we construct the test set by randomly sampling $\frac{1}{5}$ of human-annotated object referrals in the MultiScan dataset. This results in a test set with around 1.7K object referrals randomly drawn from 8.5k human-annotated object referrals in the MultiScan dataset. In the zero-shot settings, we use all scene-text pairs from datasets in SceneVerse except for ScanNet and MultiScan. This includes both human-annotated and generated texts in ARKitScenes, 3RScan, and HM3D. This setting serves to test models’ generalization capability in grounding objects with both unseen scenes and unseen texts. In the zero-shot text setting, we add generated scene-text pairs in ScanNet and MultiScan into the data used in the zero-shot setting, thereby making the held-out test containing mainly unseen object referrals.
Implementation
In the zero-shot experiments, we mainly considered three model settings scratch, zero-shot, and zero-shot text. For the zero-shot setting, we pre-train the model following Sec. B.2 without additional grounding heads considering there is no additional training data available in the zero-shot transfer setting. In the scratch and zero-shot text setting, we follow the model implementation described in Sec. C.1 and add an additional 2-layer MLP over the object features from the self-attention transformer. We follow the same fine-tuning setting described in Sec. C.1.
Appendix D Additional Results
In this section, we provide additional experimental results. Specifically, we leverage our collected SceneVerse as the pre-training data source for a traditional 3D semantic segmentation task. Next, we provide ablative analyses of our model design.
D.1 Semantic Segmentation
Setting
To test if the scaling effect of SceneVerse is universally beneficial for 3D understanding tasks, we use 3D semantic segmentation as a signature task to illustrate the effectiveness of SceneVerse. Notably, a recent work that introduced the Swin3D model [85] has identified the importance of pre-training for 3D semantic segmentation [85]. Following the same setting, we test if the proposed Swin3D model could be further improved by substituting the pre-training data to SceneVerse. Specifically, we test models’ performance on the ScanNet semantic segmentation task with 20 semantic categories and report the mean IoU and mean Acc on the validation set of ScanNet. As the original implementation of Swin3D pre-training requires surface normals as additional inputs, we reimplement the model and pre-train all models with only point coordinates and colors.
Table A.3: Semantic segmentation results on ScanNet validation set. normal-$\dagger$ denotes model trained with surface normals as an additional input. S3D indicates models initialized with the original Swin3D model weights pre-trained on Structured3D provided by Yang et al. [85].
Methods Init. SceneVerse Pre. mIoU mAcc
[…]
Swin3D-S 63.2 72.8
Swin3D-S S3D 64.1 75.1
Swin3D-S (pre-train) 67.7 78.0
Swin3D-S (pre-train) S3D 69.5 80.1
Swin3D-S (fine-tuned) S3D 70.6 80.2
Comparison
As shown in Tab. A.3, we observe a significant model performance improvement ($\sim$6%) by training Swin3D-S model on our SceneVerse dataset. Comparing our pre-training set to Structured 3D, we also observe consistent model performance improvement, showcasing the benefit of scaling-effect in SceneVerse. Moreover, we fine-tune the model on ScanNet after pre-training on SceneVerse. This process further brings improvement in model performance on semantic segmentation. We believe these results serve as strong pieces of evidence validating the effectiveness of data scaling in SceneVerse and also its potential benefit for all 3D tasks in addition to 3D visual grounding.
D.1.1 Model Ablation
In this section, we provide ablative analyses of our multi-level contrastive alignment design. We mainly consider removing objectives in our model as ablations. We choose the referral-object-level alignment objective as the default setting and consider removing: (i) object-level alignment objective, (ii) masked language modeling objective, and (iii) scene-level alignment objective. For removing the object-level alignment objective, we remove the first stage pre-training of the object point cloud encoder and jointly learn this module within the referral-object-level alignment. As shown in Tab. A.4, we test different model settings on the SceneVerse-val without additional fine-tuning. First, we show that the scene-level alignment objective is crucial for referral object grounding in SceneVerse-val with the $\sim$5performance drop. Similar observations could be made for the model trained without object-level alignment ($\sim$2drop) and masked language modeling objective ($\sim$1.5drop). We believe these results affirm the effectiveness of our overall model design.
Table A.4: Model ablation of our model on SceneVerse-val.
Obj-lvl MLM Scene-lvl Overall Easy Hard
64.8 75.4 48.7
65.2 77.1 47.4
62.4 73.4 45.8
66.9 77.8 50.3

### v3

@@
-SceneVerse: Scaling 3D Vision-Language Learning for Grounded Scene Understanding
+State Key Laboratory of General Artificial Intelligence, BIGAI %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.https://scene-verse.github.io
 Abstract
-3D vision-language grounding, which focuses on aligning language with the 3D physical environment, stands as a cornerstone in the development of embodied agents.
-In comparison to recent advancements in the 2D domain, grounding language in 3D scenes faces several significant challenges: (i) the inherent complexity of 3D scenes due to the diverse object configurations, their rich attributes, and intricate relationships; (ii) the scarcity of paired 3D vision-language data to support grounded learning; and (iii) the absence of a unified learning framework to distill knowledge from grounded 3D data.
-In this work, we aim to address these three major challenges in 3D vision-language by examining the potential of systematically upscaling 3D vision-language learning in indoor environments.
-We introduce the first million-scale 3D vision-language dataset, SceneVerse, encompassing about $68$K 3D indoor scenes and comprising $2.5$M vision-language pairs derived from both human annotations and our scalable scene-graph-based generation approach.
-We demonstrate that this scaling allows for a unified pre-training framework, Grounded Pre-training for Scenes (GPS), for 3D vision-language learning.
-Through extensive experiments, we showcase the effectiveness of GPS by achieving state-of-the-art performance on all existing 3D visual grounding benchmarks.
-The vast potential of SceneVerse and GPS is unveiled through zero-shot transfer experiments in the challenging 3D vision-language tasks.
+3D vision-language (3D-VL) grounding, which aims to align language with 3D physical environments, stands as a cornerstone in developing embodied agents.
+In comparison to recent advancements in the 2D domain, grounding language in 3D scenes faces two significant challenges: (i) the scarcity of paired 3D-VL data to support grounded learning of 3D scenes, especially considering complexities within diverse object configurations, rich attributes, and intricate relationships; and (ii) the absence of a unified learning framework to distill knowledge from grounded 3D data.
+In this work, we aim to address these major challenges in 3D-VL by examining the potential of systematically upscaling 3D-VL learning in indoor scenes.
+We introduce the first million-scale 3D-VL dataset, SceneVerse, encompassing $68$K indoor scenes and comprising $2.5$M vision-language pairs collected from both human annotations and our scalable scene-graph-based generation approach.
+We demonstrate that this scaling allows for a unified pre-training framework, Grounded Pre-training for Scenes (GPS), for 3D-VL learning.
+Through extensive experiments, we showcase the effectiveness of GPS by achieving state-of-the-art performance on existing 3D visual grounding and question-answering benchmarks.
+We also show that the data scaling effect is not limited to GPS, but is generally beneficial for models on tasks like 3D semantic segmentation.
+The vast potential of SceneVerse and GPS is unveiled through zero-shot transfer experiments in challenging 3D-VL tasks.
+Keywords:
+3D Vision-Language Data Scaling Grounded Scene Understanding
+1 Introduction
+The foundation of human cognitive development lies in the grounding of language within the physical world [84, 55, 111].
+Recent progress in Large Language Models (LLMs[11, 86, 10] has markedly promoted the alignment between vision and language [78, 3, 61] utilizing billion-scale vision-language datasets [82, 110].
+However, with these advancements predominantly focusing on the 2D domain, the grounded understanding of 3D physical environments remains in an incipient stage [16, 1, 5].
+Recognizing the pivotal role of grounded 3D experiences in shaping human cognition [7, 8], there is a compelling need to focus on exploring vision-language learning in the context of 3D scenes.
 Figure 1: Overview of SceneVerse.
 A million-scale 3D vision-language dataset that comprises over $68$K various 3D indoor scenes and $2.5$M aligned scene-language pairs in the form of scene caption, object caption, and object referral.
-$*{}^{*}$ indicates equal contribution.
-1 Introduction
-The foundation of human cognitive development lies in the grounding of language within the physical world [73, 46, 97].
-Recent progress in Large Language Models (LLMs[11, 75], often referred to as “foundation models” [10], has markedly promoted the alignment between vision and language [66, 3, 51] through utilizing billion-scale vision-language datasets [71, 96].
-Nonetheless, with these advancements predominantly focusing on the 2D domain, the grounded understanding of 3D physical environments remains in an incipient stage [16, 1, 5].
-Recognizing the pivotal role of grounded 3D experiences in shaping human cognition [7, 8] and the delayed research development in this area, there is a compelling need to intensify the exploration into the vision-language learning challenge, specifically in the context of 3D scenes.
-Seeking insights from the 2D vision-language (2D-VL) achievements, a major factor to the success was the notable scale-up of paired vision-language data [45, 15, 71].
-However, applying these principles directly from 2D to 3D is fraught with challenges.
+Seeking insights from success in 2D vision-language (2D-VL) learning, a major factor to the success was the notable scale-up of paired vision-language data [54, 15, 82].
+However, applying this experience directly from 2D to 3D is fraught with challenges.
 Primarily, 3D data collection heavily relies on the scanning device, making it inherently much more complex and expensive than gathering 2D images.
@@
 These unique aspects of 3D scenes not only make the accurate description of objects and their relations more challenging but also considerably increase the number of language descriptions required for thorough scene depiction.
-Consequently, this presents a significant challenge in obtaining a sufficient supply of high-quality paired scene-language data crucial for grounded scene understanding.
-To confront these challenges, we propose consolidating current efforts to build up SceneVerse, the first million-scale dataset aimed at advancing 3D vision-language (3D-VL) learning for grounded scene understanding.
-At the scene level, we unify 3D scene data from existing datasets [23, 78, 9, 67, 40] and supplement the collection with synthetic scenes [95, 27].
-This compilation represents the most extensive 3D scene data gathered to date, amounting to $68,406$ scenes for grounding.
-Additionally, we propose an automated generation pipeline utilizing 3D scene graphs [4, 79] and LLMs to create comprehensive, high-quality scene-language pairs.
-This refined collection, including $190,836$ human annotated pairs and totaling $2.5$M scene-language pairs, provides detailed and comprehensive portrayals of both object-level and scene-level descriptions within the 3D scene.
-We thoroughly investigate the potential offered by the data scale-up in SceneVerse with large-scale pre-training.
-Specifically, we present a novel and unified pre-training framework, Grounded Pre-training for Scenes (GPS), which is designed with scene-level and object-level alignment objectives and devoid of auxiliary losses and designs.
-Through multi-level contrastive alignment, we observe significant performance improvements across all existing 3D visual grounding benchmarks, achieving new state-of-the-art results through a simple and effective pre-training process.
-Additionally, we unveil the vast possibilities offered by SceneVerse and GPS in 3D-VL tasks in a zero-shot transfer setting.
-At last, we provide a more comprehensive understanding of the data-scaling effects in SceneVerse through extensive ablative experiments to point out future directions.
-Our main contributions can be summarized as follows:
+Consequently, this presents a significant challenge in gathering sufficient and high-quality paired scene-language data for grounded scene understanding.
+To confront these challenges, we propose SceneVerse, the first million-scale dataset aimed at advancing 3D vision-language (3D-VL) learning for grounded scene understanding.
+At the scene level, we unify 3D scene data from existing datasets [25, 88, 9, 79, 68], aligning scenes and annotations from various capturing sources, and supplement the collection with synthetic scenes [109, 29].
+This compilation represents the most extensive 3D scene data gathered to date, amounting to $68$K scenes.
+For language, we first present $97$K newly-annotated referring expressions, the most extensive thus far.
+We additionally propose an automated generation pipeline utilizing 3D scene graphs [4, 89] and LLMs to create comprehensive, high-quality scene-language pairs.
+This refined collection, totaling $2.5$M scene-language pairs, provides detailed and comprehensive descriptions of both object-level and scene-level descriptions within the 3D scene.
+We thoroughly investigate the potential offered by SceneVerse with large-scale pre-training, introducing Grounded Pre-training for Scenes (GPS), a novel and unified pre-training framework designed for scene and object-level alignment without auxiliary losses.
+Through multi-level contrastive alignment, we achieve significant performance boosts on 3D-VL tasks, such as grounding and question answering, setting new state-of-the-art results via a simple and effective pre-training process.
+We unveil the vast possibilities offered by SceneVerse and GPS in 3D-VL tasks in a zero-shot transfer setting.
+Additionally, we show that the scaling effect in SceneVerse is not limited to GPS, but generally benefits models in tasks like 3D semantic segmentation.
+At last, we offer deeper insights into the data-scaling in SceneVerse through extensive ablative experiments, pointing out future directions.
+Our main contributions are as follows:
+  1. 1.
     We introduce SceneVerse, the first million-scale 3D-VL dataset for grounded scene understanding.
@@
 This represents a significant improvement in terms of data diversity and scale compared to prior datasets.
-  • We propose GPS, an efficient transformer-based model trained with multi-level scene-text alignment that achieves state-of-the-art results on all existing 3D-VL grounding benchmarks, benefiting from pre-training on multi-level scene-language pairs in SceneVerse.
-  • We demonstrate that with the data scale-up and model design, our pre-trained models exhibit emerging zero-shot generalization capabilities in grounded scene understanding, paralleling the successes seen in 2D-VL models.
+  2. 2.
+    We propose GPS, a transformer-based model trained with multi-level scene-text alignment that achieves state-of-the-art results on existing 3D-VL grounding and question-answering benchmarks by pre-training on SceneVerse.
+  3. 3.
+    We demonstrate that with the data scale-up and model design, our pre-trained GPS exhibit emerging zero-shot generalization capabilities in grounded scene understanding.
+We also show that this scaling effect is not limited to GPS, but is generally beneficial for models on tasks like semantic segmentation.
 2 Related Work
 Datasets for Grounded 3D Understanding
-Obtaining aligned 3D-language data is an inherently difficult task.
-In 3D object modeling, pioneering works like ShapeNet [14] sourced 3D assets from online repositories, leading to follow-up proliferation of high-quality 3D object datasets [60, 22, 81].
+Obtaining aligned 3D-language data is a inherently difficult task.
+In 3D object modeling, pioneering works like ShapeNet [14] sourced 3D assets from online repositories, leading to a proliferation of high-quality 3D object datasets [71, 24, 94].
 Notably, recent developments include internet-scale data collection with Objaverse [28, 27], accompanied by the integration of object-level captions [96] for 3D-language alignment.
@@
 In contrast, developing datasets for grounded 3D scene understanding is even more challenging due to the extensive requirements for scene acquisition and annotation.
-Existing works curate RGB-D and scanned indoor scene datasets [23, 13, 78, 9, 67, 58] initially used for benchmarking classical grounding tasks like 3D object detection and segmentation [30, 42, 59, 77, 72].
-These semantically labeled scenes are subsequently applied in fine-grained scene grounding tasks like object referral [16, 1, 93], captioning [20, 88, 17, 19], vision-language-navigation [56, 80, 63, 38] and reasoning [5, 57, 37].
-Recent work exploits the representation of 3D scene graphs (3DSGs) [4, 79, 69], which concisely describes scenes with hierarchical structures.
-This representation is notably advantageous for planning [2, 68] and captioning [33], owing to its compatibility with LLMs.
-Nevertheless, as shown in Tab.
-1, these datasets are significantly constrained in both scene and language scales, underscoring the need for scaling up fine-grained scene-language-aligned data to enhance grounded scene understanding.
+Existing works curate RGB-D and scanned indoor scene datasets [25, 13, 88, 9, 79, 68] and synthetic scenes [29, 109, 52, 98] used for benchmarking tasks like 3D object detection and segmentation [32, 48, 69, 87, 83].
+These semantically labeled scenes are subsequently used in fine-grained scene grounding tasks like object referral [16, 1, 107], captioning [22, 102, 17, 19], vision-language-navigation [66, 91, 75, 41] and reasoning [5, 67, 42].
+Recent works exploit the representation of 3D scene graph (3DSG) [20, 4, 89, 81], which concisely describes scenes with hierarchical structures.
+This representation is notably advantageous for planning [2, 80] and captioning [37], owing to its compatibility with LLMs for flexible description generation [42, 43].
+Nonetheless, as shown in Tab.
+1, most datasets are constrained in both scene and language scales, underscoring the need for scaling up fine-grained and aligned scene-language data to enhance grounded scene understanding.
 Vision-Language Learning
-Recent years have witnessed tremendous progress in 2D vision-language learning [76, 66, 70, 3, 49, 51, 24], empowered by transformer-based pre-training models [28, 11, 62] and large-scale image-language datasets [15, 71].
-A central theme across language and 2D-VL domains is the effectiveness of data scaling [43], as demonstrated by improved alignment and expanded capabilities in open-vocabulary understanding [50, 47, 32, 44] through a simplified contrastive pre-training pipeline [66].
-However, in grounded scene understanding, the primary challenge for models has been the limited availability of paired 3D scene-language data, which restricts the application of insights gained from 2D-VL.
+Recent years have witnessed tremendous progress in 2D-VL [78, 3, 59, 61, 26, 21], empowered by transformer-based pre-training models [30, 11, 74] and large-scale image-language datasets [15, 82].
+A central theme across 2D-VL domains is the effectiveness of data scaling [51], as demonstrated by improved alignment and expanded capabilities in open-vocabulary understanding [60, 56, 34, 53] through a contrastive pre-training pipeline [78].
+However, in grounded scene understanding, the primary challenge for models has been the limited availability of paired 3D scene-language data, which restricts the application of insights drawn from 2D-VL.
 Current models for 3D scene grounding [39, 108, 100, 6, 64, 45, 47, 18, 95] heavily rely on task-specific knowledge in both model and loss designs or advanced optimization strategies [112].
 To bridge this gap, there has been a growing emphasis on employing pre-trained 2D-VL models for 3D-VL [105, 38, 96, 40, 106, 76, 85].
-Nonetheless, these models predominantly draw on information available from 2D-VL models (e.g., object attribute, affordance, etc.), falling short on capturing crucial information like object spatial relationships, which are only attainable through 3D data.
+Yet, these models mostly draw information available from 2D-VL models (e.g., object attribute, affordance, etc.), falling short on capturing crucial 3D information like object spatial relationships which are necessary for more fine-grained tasks such as grounded human-scene [50, 92, 46, 49, 93, 20] and robot-scene interactions modeling [72, 57, 70, 35].
 This urges the need for a multi-level alignment between language and 3D scenes, particularly regarding 3D-specific information.
-Considering the nascent stage of existing 3D pre-training methods [29, 84, 98], we believe SceneVerse and GPS have the potential to spearhead new avenues in 3D-VL research.
-Figure 2: SceneVerse collection and statistics.
-Given a 3D scene (a), our automated pipeline (c) generates three types of description including scene caption, object caption and object referral.
-(b) The comparison of different language sources and data composition.
+Considering the nascent stage of existing 3D pre-training methods [31, 97, 112, 113], we believe SceneVerse and GPS have the potential to spearhead new avenues in 3D-VL research.
 Table 1: Comparison of SceneVerse with existing 3DVL Datasets.
-SceneVerse expands the data scale of prior work by an order of magnitude.
-Anno.: human annotations.
-Syn.: template or LLM generated descriptions.
-Dataset 3D Data Language Total
-Scene Object Anno. Syn.
-ScanRefer[16] |||| |||| 52K - 52K
-ReferIt3D[1] 42K 200K 242K
-ScanQA[5] 1.5K 33K 27K - 27K
-SQA3D[57] |||| |||| - 33K 33K
-Multi3DRefer[93] 52K 10K 62K
-Cap3D[55] -  [^6] K 58K  [^6] K 724K
-ScanScribe[98] 3K 56K 94K 184K 278K
-SceneVerse 68K 1.5M 190K 2.3M 2.5M
+SceneVerse expands the data scale of prior work by order of magnitude.
+“VG” stands for Visual Grounding, “QA” for Question Answering, “PT” for Pre-training and “MT” for Multi-tasking.
+“Anno.” denotes language from human annotations and “Syn.” for template-based or LLM generated descriptions.
+Dataset 3D Task Obj. Scene Obj. Quality New Existing Total
+Scene Obj. Caption Caption Referral Check Anno. Syn. Anno. Syn.
+ScanRefer [16] |||| |||| VG 52K - - - 52K
+ReferIt3D [1] VG 42K 200K - - 242K
+ScanQA [5] 1.5K 33K QA - - - 27K - - - 27K
+SQA3D [67] |||| |||| QA - - - 33K - - - 33K
+Multi3DRefer [107] VG - 10K 52K - 62K
+Cap3D[65] -  [^6] K VG 58K  [^6] K - - 724K
+ScanScribe [112] 3K 56K PT - 90K 94K 94K 278K
+3D-LLM [42] 1.5K 186K MT - 659K - - 659K
+EmbodiedScan [90] 5K 890K VG - 970K - - 970K
+LEO [43] 3K 56K MT - 188K 235K 90K 513K
+SceneVerse 68K 1.5M VG 96K 2.1M 94K 200K 2.5M
 3 SceneVerse
-SceneVerse is the first million-scale dataset designed for grounded scene understanding.
-Our 3D scenes are curated from diverse existing datasets of both real and synthetic environments.
-Harnessing the power of 3D scene graphs and LLMs, we introduce an automated pipeline to generate comprehensive and high-quality language for both object-level and scene-level descriptions.
-We additionally incorporate the most extensive human-annotated object referrals to date, providing new training sources and benchmarks in this field.
+SceneVerse is designed for grounded scene understanding with 3D scenes curated from diverse existing datasets of both real and synthetic environments.
+Regarding language, we employ both human annotation and a novel automated generation pipeline to collect comprehensive and high-quality language for both object-level and scene-level descriptions.
+We provide details regarding data collection in the following sections.
 3.1 Scene Curation
@@
 To ensure cohesion across various sources, we conduct preprocessing steps such as room segmentation, point subsampling, axis alignment, normalization, and semantic label alignment.
-Each scan is represented by a point cloud $\mathrm{P}\in\mathbb{R}^{N\times 8}$, wherein each point is defined by its 3D coordinates, RGB color, instance id and semantic label.
+Each scan is represented by a point cloud $\mathrm{P}\in\mathbb{R}^{N\times 8}$, wherein each point is defined by its 3D coordinates, RGB color, instance id, and semantic label.
 In total, we curate $68,406$ 3D scenes in SceneVerse.
+3.2 Referral Annotation by Humans
+In the curated scenes of SceneVerse, we present the most comprehensive set of human-annotated, context-rich object referrals to date, serving as a valuable benchmark for assessing grounded scene understanding capabilities.
+The human annotations contain $96,863$ descriptions in ARKitScenes [9], HM3D [79] and MultiScan [68].
+During the annotation process, one human annotator was assigned to write at least 20 words to distinctly refer to a single 3D object within a 3D scene.
+Each referral text then undergoes independent verification by two additional reviewers, both mandated to accurately locate the referenced object based on the 3D scene and the annotated referral text.
+Any object referrals that do not pass the verification by either reviewer are flagged for re-annotation.
+Figure 2: SceneVerse collection and statistics.
+Given a 3D scene (a), our automated pipeline (c) generates three types of description including scene caption, object caption and object referral.
+(b) SceneVerse data comparison and composition.
 3.3 3D Scene Graph Construction
-Our 3D scene graph is defined as a set of tuples $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the nodes $\mathcal{V}$ comprises $\mathcal{V}_{1}\bigcup\mathcal{V}_{2}\bigcup\dots\bigcup\mathcal{V}_{K}$, with $\mathcal{V}_{k}$ representing the set of nodes at a particular hierarchical level.
-Each node $v$ represents one distinct 3D object instance, parameterized by its centroid ${\bm{p}}_{i}\in\mathbb{R}^{3}$ and bounding box size of ${\bm{b}}_{i}=(b_{x},b_{y},b_{z})\in\mathbb{R}^{3}$.
+Our 3D scene graph is defined as a hierarchical graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
+Each node $v\in\mathcal{V}$ represents one distinct 3D object instance, parameterized by its centroid ${\bm{p}}_{i}\in\mathbb{R}^{3}$ and bounding box size of ${\bm{b}}_{i}=(b_{x},b_{y},b_{z})\in\mathbb{R}^{3}$.
 The edges $\mathcal{E}$ represent spatial relationships between nodes.
-To construct the scene graph $\mathcal{G}$, we first instantiate the nodes with the instance annotation from the point clouds and assign object classes with their corresponding semantic labels.
-Following prior work[1, 79], we consider the following spatial relations.
-Vertical proximity
-This encompasses both in-contact relationships (e.g., support, inside, embed), and non-contact ones (e.g., above, below).
-Horizontal proximity
-Horizontal relationships describe the proximity relations like in front of, next to, behind, etc.
-Relationships like left, right are contextually dependent on a reference view, where another anchor object is utilized to establish the view direction.
-The distance between the two objects is also calculated to describe whether the objects are far or near in space.
-Multi-object Relationships
-This models the spatial arrangement of multiple objects, e.g., align and between.
-The node hierarchy is decided by the support relationship.
-We traverse all the object nodes to calculate spatial relationships, which undergo an automatic verification procedure to rectify incorrect ones.
-For a more detailed description of the scene graph construction and relationship determination, please refer to Sec.
-A.2.
+To construct the scene graph $\mathcal{G}$, we instantiate the nodes with the instance annotation from the point clouds and assign object classes with their corresponding semantic labels.
+Following prior work[1, 89], we consider the Vertical proximity, Horizontal proximity and Multi-object Relationships as spatial relations.
+For a more detailed description of the scene graph construction and relationship determination, please refer to supplementary.
 3.4 Language Generation with LLMs
-The scene-language pairs in SceneVerse aim to capture varying aspects of the 3D scene, which include detailed object attribute descriptions in object captioning, spatial relationships between objects in object referral, and global scene descriptions in scene captioning.
-Based on the 3D scene graph, we utilize both templates and LLMs to automatically generate descriptions on these three granularities.
+The scene-language pairs in SceneVerse aim to capture varying aspects of the 3D scene, including detailed object attributes in object captioning, spatial relationships between objects in object referral, and global scene descriptions in scene captioning.
+Based on 3D scene graphs, we utilize both templates and LLMs to automatically generate descriptions on these three granularities.
 Object Captioning
@@
 The images are then cropped with the rendered bounding boxes and processed through BLIP2 [58] to generate initial object captions.
-To refine the captions, we select the top 10 sentences with the highest CLIP [66] similarity score and minimal occlusion.
-The selected sentences are fed into a LLM to obtain a coherent summary of the object captions.
-In this process, we explicitly instruct the language model to identify and correct the potential errors.
-The detailed object captioning pipeline is illustrated in Sec.
-A.3.
+We select the top 10 sentences with the highest CLIP [78] similarity score and minimal occlusion and utilize an LLM to obtain a refined and coherent summary of the captions.
+The detailed object captioning pipeline is illustrated in supplementary.
 Object Referral
@@
 The scene-level captions emphasize global information, portraying the key objects in the scene along with their attributes and functionalities.
-We leverage the constructed 3D scene graph and prompt LLMs to generate these captions.
-To enhance the diversity of scene captions, we utilize a subgraph sampling strategy, where a subset of edges and nodes are randomly sampled as the scene context.
+We use the constructed 3D scene graph and prompt LLMs to generate these captions.
+We random sample a subset of edges and nodes from the scene graph each time as the scene context to enhance the diversity of scene captions.
 The object counts are also provided as LLM prompts, together with the room type and object attributes if such annotations are available in the dataset.
-3.4 Referral Annotation by Humans
-In addition to automatically generated scene-text pairs, SceneVerse includes the most comprehensive set of human-annotated, context-rich object referrals to date, serving as a valuable benchmark for assessing grounded scene understanding capabilities.
-The human annotations contain $96,863$ descriptions in ARKitScenes [9], HM3D [67] and MultiScan [58].
-During the annotation process, one human annotator was assigned to write at least 20 words to distinctly refer to a single 3D object within a 3D scene.
-Each referral text then undergoes independent verification by two additional reviewers, both mandated to accurately locate the referenced object based on the 3D scene and the annotated referral text.
-Any object referrals that do not pass the verification by either reviewer are flagged for re-annotation.
-3.5 Details and Statistics
+3.5 Data Quality and Statistics
+Data Quality
+The $96$K human-annotated set of SceneVerse is collected through AMT, where $82$ humans are employed for annotation and $18$ for verification.
+All final annotations passed the reference verification, with a re-annotation rate of $4.8\%$.
+For our automatic language generation pipeline, we conduct extensive prompt tuning and iterate with human feedback for LLMs on object captioning, summary, and rephrasing.
+To verify the efficacy of the pipeline, we conduct a quality check where $12$K generated object-level descriptions are randomly selected for human verification.
+Results demonstrate a $96.93\%$ pass rate, surpassing that in ReferIt3D [1] with $86.1\%$ pass rate on $2$K samples.
+Statistics
 In total, SceneVerse comprises a total of $68,406$ room-level 3D scans, with the source composition shown in Fig.
 2 (b).
-The dataset contains $1.5$M object instances, comprising $21$ types of relationships following prior work [1, 79].
-For the language descriptions, we generate $1$M template-based texts and $1$M sentences by LLM rephrased by Llama [75] and GPT-3.5 [61].
-All the rephrasing and summary prompts, along with the complete set of relationships, are detailed in Sec.
-A.3.
-To verify the efficacy of our automated language generation pipeline, we conduct a quality check (QC) where 12K generated object-level descriptions are randomly selected for human verification, achieving a $96.93\%$ pass rate.
-This shows the capability of our proposed scene-graph-based generation approach to produce high-quality language descriptions, laying a robust foundation for future scalability.
+The dataset contains $1.5$M object instances ranging in 2290 object categories.
+Our generated 3D scene graph comprises $21$ types of relationships following prior work [1, 89].
+For the language descriptions, we generate $1$M template-based texts and $1$M sentences rephrased by Llama [86] and GPT-3.5 [73].
+As can be seen from the radar chart and examples in Fig.
+2, the diversity of the LLM-refined data, particularly in sentence length and variety, closely aligns with the characteristics of annotated descriptions, surpassing the template-based data.
+Together with existing sources (294K) and our newly annotated set (96K), SceneVerse contains $2.5$M scene-language pairs in total.
+All the rephrasing and summary prompts, along with the complete set of relationships, are detailed in supplementary.
 4 Grounded Pre-training for Scenes
-Figure 3: Overview of our proposed GPS model.
-We leverage contrastive alignment in three levels $\mathcal{L}_{\text{obj}}$, $\mathcal{L}_{\text{scene}}$, and $\mathcal{L}_{\text{ref}}$ as well as a masked language modeling objective $\mathcal{L}_{\text{MLM}}$ for model learning.
+Figure 3: Overview of GPS model.
+We use contrastive alignment at three levels $\mathcal{L}_{\text{obj}}$, $\mathcal{L}_{\text{scene}}$, and $\mathcal{L}_{\text{ref}}$ and a masked language modeling objective $\mathcal{L}_{\text{MLM}}$ for model learning.
 In this section, we introduce GPS, an efficient transformer-based model trained with multi-level contrastive losses for aligning 3D scenes and texts.
 As shown in Fig.
-3, we echo the language descriptions collected at different levels to form scene-language pairs at both object-level, referral-object-level, and scene-level for contrastive objectives in GPS.
+3, we echo the language descriptions collected at different granularities to form contrastive objectives at both object-level, referral-object-level, and scene-level in GPS.
 We describe the design of each level in the following sections.
@@
 Following [96], we perform cross-modal alignment on the object features and text features via:
-$\mathcal{L}_{\text{obj}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(p,r)\right)}}+\right.$ (1)
-$\left.\ \ \ \log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(r,q)\right)}}\right),$
+$\mathcal{L}_{\text{obj}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(p,r)\right)}}+\log\frac{\exp{\left(D^{\text{obj}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{obj}}(r,q)\right)}}\right),$ (1)
 where $D^{\text{obj}}(p,q)=({\bm{f}}^{O}_{p}{\bm{f}}^{T}_{q}/\tau)$ denotes the dot product between object and text features and $(p,q)$ denotes a pair of aligned object-text pair in the training batch and $r$ iterates over all object-text pairs in the training batch.
@@
 Similar to object-level grounding, we pass the scene caption through a tunable language model to obtain text feature ${\bm{g}}^{T}$ and perform scene-level contrastive alignment through:
-$\mathcal{L}_{\text{scene}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(p,r)\right)}}+\right.$ (2)
-$\left.\ \ \ \log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(r,q)\right)}}\right),$
+$\mathcal{L}_{\text{scene}}=-\frac{1}{2}\sum_{(p,q)}$ $\left(\log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(p,r)\right)}}+\log\frac{\exp{\left(D^{\text{scene}}(p,q)\right)}}{\sum_{r}\exp{\left(D^{\text{scene}}(r,q)\right)}}\right),$ (2)
 where $D^{\text{scene}}(p,q)=({\bm{g}}_{p}^{S}{\bm{g}}_{q}^{T}/\tau)$ denotes the dot product between scene feature ${\bm{g}}_{p}^{S}$ and scene caption feature ${\bm{g}}_{q}^{T}$ for each pair of aligned scene-text pairs in the training batch and $r$ iterates over all scene-text pairs in the training batch.
@@
 This mimics the success of intra-image and inter-image contrasts commonly used for region-word alignment in 2D-VL models [104].
-To learn the multi-level alignment between 3D scenes and language, we first train the point cloud encoder with object-level grounding objects to obtain a good feature initialization for grounding objects in scenes.
-During the scene grounding stage, we train our inter- and intra-scene objectives together with a mask language modeling loss $\mathcal{L}_{\text{MLM}}$ over the inputted object-referral texts to tune the parameters within the language encoder and self-attention transformer.
-Above all, the learning of GPS could be summarized as optimizing the following objective:
+To learn the multi-level alignment between 3D scenes and language, we first train the point cloud encoder with an object-level grounding objective to obtain a good feature initialization for grounding objects in scenes.
+During the scene grounding stage, we train our inter- and intra-scene objectives together with a masked language modeling loss $\mathcal{L}_{\text{MLM}}$ over the inputted object-referral texts to tune the parameters within the language encoder and self-attention transformer.
+Above all, the learning of GPS could be summarized as optimizing:
 $\mathcal{L}=\mathcal{L}_{\text{obj}}+\mathcal{L}_{\text{scene}}+\mathcal{L}_{\text{ref}}+\mathcal{L}_{\text{MLM}}.$
 5 Experiments
-In this section, we present the evaluation results addressing the following questions:
-  • How effective is the data scaling in SceneVerse for 3D visual grounding? Does the scale-up work for general pre-training based 3D-VL models?
-  • How well is the GPS pre-training pipeline? Does it exhibit similar properties of 2D-VL models in 3D-VL tasks?
-  • What potentials are offered by SceneVerse and GPS for future research? What is missing?
-In the following sections, we describe in detail about model performance regarding these key topics.
-Due to the page limit, we direct readers to the Appendices B and C for implementation details and more experimental analyses.
+In this section, we present experimental results addressing the following questions:
+  1. 1.
+    How effective is the data scaling in SceneVerse for 3D visual grounding?
+Does the scale-up benefit common 3D-VL tasks (e.g., 3D question answering, open-vocabulary 3D semantic segmentation) and pre-training-based models?
+  2. 2.
+    How well is the GPS pre-training pipeline for 3D-VL tasks? Does it exhibit similar properties of 2D-VL models in 3D-VL tasks?
+  3. 3.
+    What is offered by SceneVerse and GPS and what is missing?
+In the following sections, we describe in detail the model performance regarding these key topics.
+Due to the page limit, we direct readers to the supplementary for implementation details, qualitative results, and more experimental analyses.
 5.1 3D Visual Grounding
+Settings
+We evaluate our model on three commonly-used datasets for 3D visual grounding: ScanRefer [16], Nr3D, and Sr3D [1].
+For Nr3D and Sr3D, we follow Achlioptas et al.
+[1] and report the grounding accuracies of models using ground-truth object masks.
+For ScanRefer, we follow Zhu et al.
+[112] and use Mask3D [83] to generate object proposals.
+Results are reported as $\text{Acc}@0.5$ to evaluate the correctness of predictions whose object bounding boxes overlap the ground truth with IoU $>0.5$.
+For comparisons, we compare with existing baselines by providing the results of pre-trained GPS and dataset-specific fine-tuned GPS.
+Please see more details in the supplementary.
+Results and Analyses
+As shown in Tab.
+2, GPS trained on SceneVerse achieves state-of-the-art results on all existing 3D-VL grounding benchmarks.
+Initially, when GPS is trained directly on the training sets of benchmark datasets, labeled as Ours (scratch), it underperforms compared to existing models that employ more complex structures or loss designs.
+This result underscores the data-intensive nature of the contrastive alignment paradigm.
+However, when presented with extensive training data in SceneVerse, the results of our model without additional fine-tuning, i.e., Ours (pre-train), significantly improves and already achieves state-of-the-art results on benchmarks like ScanRefer.
+Moreover, the dataset-specific fine-tuned model, i.e., Ours (fine-tuned), consistently outperforms existing baselines with only a simple projection MLP added on top of the pre-trained model, jointly optimized during fine-tuning without any other auxiliary architecture or loss objective.
+These results underscore the strong potential of both the SceneVerse and GPS for 3D-VL tasks.
 Table 2: 3D visual grounding results on Nr3D, Sr3D, and ScanRefer.
-We use “direct” for our model trained on SceneVerse with no additional fine-tune head, and “fine-tune” for the data-specific fine-tuned version of our model.
-We highlight the best results in bold.
+We use “pre-train” for our model trained on SceneVerse w/o additional fine-tuning, and “fine-tune” for its data-specific fine-tuned version.
+Best results are highlighted in bold.
 Method Nr3D Sr3D ScanRefer Acc@0.5
@@
 Ours (fine-tuned) 64.9 72.5 57.8 56.9 67.9 77.5 80.1 71.6 62.8 78.2 48.1 77.9 42.7
-Settings
-We evaluate our model on three commonly-used datasets for 3D visual grounding: ScanRefer [16], Nr3D, and Sr3D [1].
-For Nr3D and Sr3D, we follow Achlioptas et al.
-[1] and report the grounding accuracies of models using ground-truth object masks.
-For ScanRefer, we follow Zhu et al.
-[98] and use Mask3D [72] to generate object proposals.
-Results are reported as $\text{Acc}@0.5$ to evaluate the correctness of predictions whose object bounding boxes overlap the ground truth with IoU $>0.5$.
-For comparisons, we compare with existing baselines by providing the results of pre-trained GPS and dataset-specific fine-tuned GPS.
-Please see more details in the Appendix C.
-Results and Analyses
-As shown in Tab.
-2, GPS trained on SceneVerse achieves state-of-the-art results on all existing 3D-VL grounding benchmarks.
-Initially, when GPS is trained directly on the training sets of benchmark datasets, labeled as Ours (scratch), it underperforms compared to existing models that employ more complex structures or loss designs.
-This result underscores the data-intensive nature of the contrastive alignment paradigm.
-However, when presented with extensive training data in SceneVerse, the results of our model without additional fine-tuning, i.e., Ours (pre-train), significantly improves and already achieves state-of-the-art results on benchmarks like ScanRefer.
-Moreover, the dataset-specific fine-tuned model, i.e., Ours (fine-tuned), consistently outperforms existing baselines with only a simple projection MLP added on top of the pre-trained model, jointly optimized during fine-tuning without any other auxiliary architecture or loss objective.
-These results underscore the strong potential of both the SceneVerse and GPS for 3D-VL tasks.
 5.2 Zero-Shot Transfer
@@
 We create SceneVerse-val using $8.5K$ annotated object referrals of $271$ scenes in MultiScan, and randomly split the scenes following a 4:1 train / test split for creating the held-out test set.
-We mainly consider 2 specific transfer settings in our experiments: (i) zero-shot: models trained by removing all the scenes from the target dataset, tested on held-out unseen scenes, and (ii) zero-shot text: Models trained on data that include the 3D scenes from training set of the target dataset, yet tested exclusively with unseen scene-text distribution.
+We mainly consider 2 specific transfer settings in our experiments: (i) zero-shot: models trained by removing all the scenes from the target dataset, tested on held-out unseen scenes, and (ii) zero-shot text: Models trained on data that include the training set of scenes from the target dataset, yet tested exclusively with unseen scene-text distribution.
 Specifically, for the zero-shot text setting, we use the generated texts in SceneVerse as fine-tuning sources for the zero-shot model.
 We mainly compare our model against a recent pre-training-based model 3D-VisTA.
-See more details on experimental setting and implementation in the Appendix C.
-Table 3: Zero-shot transfer results on established benchmarks.
-Method Nr3D Sr3D ScanRefer@0.25 ScanRefer@0.5
+See more details on experimental setting and implementation in the supplementary.
+Table 3: Zero-shot transfer on existing benchmarks. “SR” stands for ScanRefer.
+Method Nr3D Sr3D SR@0.25 SR@0.5
 3D-VisTA (scratch) 57.5 69.6 45.9 41.5
@@
 Ours (zero-shot text) 41.9 38.1 40.7 35.8
-Table 4: Zero-shot transfer on SceneVerse-val. We evaluate models following settings in Nr3D/Sr3D using GT object proposals.
+Table 4: Zero-shot transfer on SceneVerse-val. Evaluation uses GT object proposals following Nr3D/Sr3D.
 Method Overall Easy Hard V-Dep. V-Indep.
@@
 Results and Analyses
-We present the results of zero-shot transfer experiments in Tab. 3 and Tab. 4 with the following key observations:
+We present the results of zero-shot transfer experiments in Sec. 5.2 and Sec. 5.2 with the following key observations:
+[…]
     Our GPS model demonstrates superior generalization to unseen scenes compared to the 3D-VisTA model.
@@
 This indicates the effectiveness of contrastive alignment over traditional classification objectives, aligning with the advancements seen in 2D-VL models for open-vocabulary grounding and transfer capabilities
+[…]
     SceneVerse dataset substantially enhances 3D-VL grounding capabilities through zero-shot transfer, especially when provided with relatively limited training data, i.e., SceneVerse-val.
-As demonstrated in  Tab.
-4, there is a significantly improved performance when comparing models trained on SceneVerse in a zero-shot manner to those trained from scratch.
+As demonstrated in  Sec.
+5.2, there is a significantly improved performance when comparing models trained on SceneVerse in a zero-shot manner to those trained from scratch.
 This indicates that SceneVerse can effectively capture knowledge for general 3D scene grounding.
 Consequently, this underscores its potential as a go-to pre-training dataset for 3D-VL tasks.
+[…]
     The impact of our extensive collection and scalable generation of scene-text pairs is further evidenced by the results in the zero-shot text setting.
-Notably, as shown in Tab.
-3, these automatically generated scene-text pairs supply ample knowledge for comprehending the scene distribution.
+Notably, as shown in Sec.
+5.2, these automatically generated scene-text pairs supply ample knowledge for comprehending the scene distribution.
 This contributes significantly to the substantial improvement over the zero-shot performance.
-Figure 4: Model performance v.s. data scale. Models consistently improve in both the pre-train and zero-shot transfer settings on ScanRefer and SceneVerse-val with data scaling-up.
-5.3 Ablative Studies and Discussions
-In this section, we present ablative studies focused primarily on the data collected in SceneVerse.
-Our aim is to elucidate the effects of data scaling more clearly.
-For discussions regarding ablation studies on model architecture, readers are referred to the Appendix D.
-The following points are specifically discussed in this section.
+Table 5: 3D question answering results on ScanQA and SQA3D. We report EM@1 score on ScanQA and SQA3D evaluation sets.
+Model ScanQA SQA3D
+val w/obj w/o obj
+ScanRefer+MCAN [5] 18.6 20.6 19.0 -
+ScanQA [5] 20.3 23.5 20.9 46.6
+SQA3D [67] - - - 47.2
+3D-VisTA [112] 22.4 27.0 23.0 48.5
+3D-LLM [42] 20.5 19.1 - -
+Ours 22.7 25.0 23.5 49.9
+Table 6: Exisiting 3D backbones pre-trained on SceneVerse for open-vocabulary 3D semantic segmentation on ScanNet. “SPUNet” denotes SparseUNet proposed in [97].
+Model Network mIoU $\Delta$ mAcc $\Delta$
+OpenScene [76] SPUNet16 57.2 - 69.9 -
+PLA [31] SPUNet16 17.7 - 33.5 -
+RegionPLC [97] SPUNet16 56.9 - 75.6 -
+RegionPLC+SceneVerse SPUNet16 58.2 +1.7% 77.3 +2.2%
+OpenScene [76] SPUNet32 57.8 - 70.3 -
+PLA [31] SPUNet32 19.1 - 41.5 -
+RegionPLC [97] SPUNet32 59.6 - 77.5 -
+RegionPLC+SceneVerse SPUNet32 61.0 +2.3% 79.7 +2.8%
+5.3 Additional 3D-VL Tasks
+Settings
+We evaluate the effectiveness of GPS and SceneVerse on additional 3D-VL tasks: (i) 3D question answering (3D-QA) on ScanQA [5] and SQA3D [67], and (ii) open-vocabulary 3D semantic segmentation (OV-Seg) on ScanNet.
+[…]
+    In the 3D-QA task, we follow Zhu et al.
+[112] and evaluate models over the exact match metic (EM@1) on the validation and test sets of ScanQA, as well as the test set of SQA3D.
+We pre-train GPS on SceneVerse and fine-tune the model on the 3D-QA dataset to compare with state-of-the-art models.
+[…]
+    In the OV-Seg task, as GPS builds upon an object-centric design and thus is not directly applicable to semantic segmentation, we consider testing the effectiveness of SceneVerse on improving existing 3D models.
+Specifically, we follow the open-vocabulary semantic segmentation settings proposed by Yang et al.
+[97] and report the mIoU and mAcc score.
+We compare with existing works by pre-training the RegionPLC [97] model on SceneVerse.
+Results and Analyses
+We present the results of 3D-QA experiments in Sec. 5.2 and the results of OV-Seg experiments in Sec. 5.2. The analyses are as follows:
+[…]
+    As shown in Sec.
+5.2, our model achieves state-of-the-art results on both benchmarks, outperforming recent strong pre-training-based baselines like 3D-VisTA and 3D-LLM.
+As SceneVerse currently contains only descriptions of objects and scenes, we believe involving more types of language descriptions (e.g., question-answer pairs, dialogues) is a promising direction for further improving model performance on these downstream tasks.
+[…]
+    As shown in Sec.
+5.2, we observe consistent performance improvement of existing 3D backbone models on this task when pre-trained with SceneVerse data.
+This result validates that the collected data in SceneVerse can effectively boost the performance of existing models on scene understanding tasks.
+We further provide results of state-of-the-art 3D models that are pre-trained on SceneVerse on the close-vocabulary 3D semantic segmentation task in the supplementary.
+5.4 Ablative Studies and Discussion
+In this section, we present further discussions on both the data collected in SceneVerse and the GPS model design.
+We aim to elucidate the effects of data scaling and show more clearly its effectiveness in 3D scene understanding.
+Regarding the experimental settings and more results discussion, refer to the supplementary.
+The following points are specifically discussed in this section:
 How important is data-scaling?
@@
 4, we observe consistent performance improvement over the increase of data scale for both settings.
-We provide additional experiments in the Appendix D to show that such scaling effect is not only beneficial for 3D-VL grounding but also for other 3D tasks like semantic segmentation [72, 85].
+We provide additional experiments in the supplementary to show that such scaling effect is not only beneficial for 3D-VL grounding but also for other 3D tasks like semantic segmentation [83, 99].
+[Uncaptioned image][Uncaptioned image]
+Figure 4: Model performance v.s. data scale. Plots show that models consistently improve in both the pre-train and zero-shot transfer settings on ScanRefer and SceneVerse-val with data scaling-up.
+Table 7: Ablation on text data source used in model pre-training. All models are tested on ScanRefer with no additional finetuning.
+Template LLM Anno. Acc@0.25 Acc@0.5
+43.5 38.4
+50.9 46.1
+51.1 46.3
+52.0 47.1
+Table 8: Cross domain transfer results of models pre-trained on real and synthetic datasets. “S3D” stands for Structured3D.
+Real Synthetic SceneVerse-val S3D ProcTHOR
+All 64.8 37.1 43.4
+S3D 7.0 85.1 16.1
+ProcTHOR 4.2 16.3 91.0
+Table 9: Ablation on model design on SceneVerse-val. We use “Obj-lvl”, “Scene-lvl” to denote object and scene alignment loss, and “MLM” for the mask language modeling loss.
+Obj-lvl. MLM Scene-lvl. Overall Easy Hard
+64.8 75.4 48.7
+65.2 77.1 47.4
+62.4 73.4 45.8
+66.9 77.8 50.3
 How is the generated data compared with human-annotated data?
 We assess the performance of models trained using various scene-text sources, specifically focusing on their performance in the ScanRefer dataset without additional fine-tuning.
-As shown in Tab.
-5, models trained with our template-based generated texts and Large Language Model (LLM)-refined texts show significant improvements over models trained solely on ScanRefer.
+As shown in Sec.
+5.4, models trained with our template-based generated texts and Large Language Model (LLM)-refined texts show significant improvements over models trained solely on ScanRefer.
 More importantly, these variants of our model already achieve state-of-the-art results compared with previous baselines.
@@
 Specifically, we compare models trained on all real scenes in SceneVerse against models trained exclusively on two synthetic subsets of SceneVerse, i.e., Structured3D and ProcTHOR.
-As shown in Tab.
-6, models trained on synthetic subsets demonstrate remarkable performance on their corresponding test sets while suffering when transferred to real or other synthetic scenes.
-In contrast, the model trained on real scene-text pairs exhibits less severe performance drops when generalizing to synthetic scenes.
-This result affirms the domain gap between real and synthetic scenes in 3D-VL grounding and shows that a simple scale-up in the amount of scenes is insufficient when the scene naturalness can not be guaranteed.
-Considering the scalability of our quality-ensured language generation and also the scaling effect shown in our experiments, the rate-determining step for further scaling-up 3D-VL comes to the collection of diverse, high-quality, and realistic scenes that capture natural 3D scene distributions.
-Table 5: Ablation on scene-text pair types used in training. We report model results on ScanRefer with no additional finetuning.
-Template LLM Anno. Acc@0.25 Acc@0.5
-43.5 38.4
-50.9 46.1
-51.1 46.3
-52.0 47.1
-Table 6: Cross domain transfer results of models learned in real and synthetic datasets without additional fine-tuning. “S3D” stands for Structured3D.
-Real Synthetic SceneVerse-val S3D ProcTHOR
-All 64.8 37.1 43.4
-S3D 7.0 85.1 16.1
-ProcTHOR 4.2 16.3 91.0
+As shown in Sec.
+5.4, models trained on synthetic subsets demonstrate remarkable performance on their corresponding test sets while suffering when transferred to real or other synthetic scenes.
+In contrast, the model trained on real scene-text pairs exhibits less performance drop when generalizing to synthetic scenes.
+This result affirms the domain gap between real and synthetic scenes in 3D-VL grounding and shows that a simple scale-up in the number of scenes is insufficient when naturalness can not be guaranteed.
+Considering the scalability of our language generation pipeline and the scaling effect shown in our experiments, the rate-determining step for further scaling-up 3D-VL comes to the collection of diverse, high-quality, and realistic scenes that capture natural 3D scene distributions.
+How important is the design of each module in GPS?
+We provide ablative analyses of our multi-level contrastive alignment design in Sec.
+5.4.
+We mainly consider removing objectives in our model to reveal the effectiveness of each level of alignment.
+We choose the referral-object-level alignment objective as the default setting and consider removing: (i) object-level alignment objective, (ii) masked language modeling objective, and (iii) scene-level alignment objective.
+When removing the object-level alignment objective, we learn the object point cloud encoder with the referral-object-level alignment and without pre-training.
+As shown in Sec.
+5.4, we test different models on the SceneVerse-val without additional fine-tuning.
+Results show that the scene-level alignment objective is crucial for referral object grounding in SceneVerse-val with the $\sim$5performance drop.
+Similar observations could be made for the model trained without object-level alignment ($\sim$2drop) and masked language modeling objective ($\sim$1.5drop).
+These results affirm the effectiveness of our model design.
 6 Conclusion
-In this work, we scale up 3D-VL in the context of grounded scene understanding.
-We introduce SceneVerse, a million-scale 3D-VL dataset encompassing various scenes and multi-level scene descriptions sourced from both human annotation and our proposed scene-text generation approach.
-Utilizing SceneVerse, we propose Grounded Pre-training for Scenes, a model trained with multi-level scene-language contrastive alignment over the data collected.
-Through extensive experiments, we show that GPS achieves state-of-the-art results on all existing 3D-VL grounding tasks.
+In this work, we scale up 3D-VL for grounded scene understanding.
+We present SceneVerse, a million-scale 3D-VL dataset covering various scenes and multi-level scene descriptions sourced from both human annotation and our proposed scene-text generation approach.
+Utilizing SceneVerse, we propose Grounded Pre-training for Scenes (GPS), a model trained with multi-level scene-language contrastive alignment.
+Through extensive experiments, we show that GPS achieves state-of-the-art results on common 3D-VL tasks including grounding and question answering.
 We further conduct zero-shot transfer experiments to show the improved generalization performances of GPS trained on SceneVerse compared with previous baselines.
-We hope our efforts and successful scale-up attempts in SceneVerse could pave the way for a new research paradigm in 3D-VL.
-7 Acknowledgement
-The authors thank Yaowei Zhang from BIGAI for designing the result visualization framework, Jiangyong Huang and Xiongkun Linghu from BIGAI for suggestions on data generation and refinement, and colleagues from BIGAI for their helpful discussions and suggestions.
+We also demonstrate that the scaling effect of SceneVerse is generally beneficial for existing 3D models on 3D-VL tasks like semantic segmentation.
+We hope our efforts and successful scale-up attempts in SceneVerse could pave the way for new research paradigms in 3D-VL.
+Acknowledgement
+The authors would like to thank Yaowei Zhang (BIGAI) for his help on online visualization and other colleagues from BIGAI General Vision Lab for fruitful discussions.
+The authors would also like to thank the anonymous reviewers for their constructive feedback.
 Supplementary Material
-In Appendix A, we introduce more details of SceneVerse, including the 3D scene preprocessing, scene graph construction and automatic language generation.
-Appendix B presents more model and implementation details.
-Appendix C include a more in-depth summary of the settings and implementations for the experiments in the main paper, as well as the ablative study and additional experiments on semantic segmentation to demonstrate the benefits of SceneVerse.
-Appendix A SceneVerse
-A.1 3D Scenes
+Appendix 0.A The SceneVerse Dataset
+0.A.1 3D Scenes
 To address the scarcity of available 3D scene data, we construct SceneVerse by unifying 3D scene data from various existing datasets.
@@
 Point Cloud Normalization
-In order to mitigate the data disparities arising from diverse capture devices across various data sources, we subsample each point cloud to a maximum of $240,000$ points.
+To mitigate the data disparities arising from diverse capture devices across various data sources, we subsample each point cloud to a maximum of $240,000$ points.
 Each point cloud then undergoes a transformation centered on the central point on the floor, followed by rotation to align the room layout with the axis following the approach by Chen et al.
@@
 In total, we curate $68,406$ 3D scenes in SceneVerse.
-A.2 3D Scene Graph Construction
+0.A.2 3D Scene Graph Construction
 In Sec. 3.3, we introduce an automated pipeline to construct 3D scene graphs from point clouds. Here, we provide more implementation details and the relationship definition.
-A.2.1 Relationships
+Relationships
 Table A.1: Relationships in SceneVerse. The 3D scene graph captures 21 types of relationships ranging in 4 categories.
 Category Relation
-|||| supported by
-In-contact vertical embedded into
-placed in
-|||| inside
-hanging on
-affixed on
-mounted on
-Non-contact vertical above
-higher than
-below
+In-contact vertical supported by embedded into
+placed in inside
+Non-contact vertical hanging on affixed on
+mounted on above
+higher than below
 lower than
-near(far) to the left of
-near(far) to the right of
-is behind
-Horizontal is in front of
-close to
-adjacent to
-besides
-next to
-Multi-object between
-aligned
+Horizontal near(far) to the left of near(far) to the right of
+is behind is in front of
+close to adjacent to
+besides next to
+Multi-object between aligned
 Our 3D scene graph captures $21$ types of relations as shown in Tab. A.1. We provide illustrations of how these relations are defined in the 3D space, as can be seen in  Fig. A.1.
-A.2.2 Scene Graph Construction
+Scene Graph Construction
 Due to the inherent noise and incompleteness in the point cloud representation, automatically extracting precise and comprehensive relationships from the point clouds is a non-trivial task.
-Below we detail our 3D scene graph construction process, as outlined in Alg.
-1.
+Below we detail our 3D scene graph construction process, as outlined in Algorithm 1.
 We first instantiate the graph nodes with the instance annotation from the point cloud and parameterize each node with object centroid $\mathrm{p_{i}}\in\mathbb{R}^{3}$ and size of the axis-aligned bounding box $\mathrm{b_{i}}=(b_{x},b_{y},b_{z})\in\mathbb{R}^{3}$ (Line 1-3).
@@
 Therefore, we would automatically refined ( TV, hanging on,  wall) to ( TV, mounted on,  wall).
-In our constructed 3D scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the nodes $\mathcal{V}$ comprises $\mathcal{V}_{1}\bigcup\mathcal{V}_{2}\bigcup\dots\bigcup\mathcal{V}_{K}$, with $\mathcal{V}_{k}$ representing the set of nodes at a particular hierarchical level.
+In our constructed 3D scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the nodes $\mathcal{V}$ comprises the union of node sets $\mathcal{V}_{1}\bigcup\mathcal{V}_{2}\bigcup\dots\bigcup\mathcal{V}_{K}$, with $\mathcal{V}_{k}$ representing the set of nodes at a particular hierarchical level.
 The hierarchies are determined by the support relationship; for instance, objects supported by the floor constitute $\mathcal{V}_{0}$, while objects supported by the table will form $\mathcal{V}_{1}$, etc.
@@
 Output : 3D scene graph $\mathcal{G}(\mathcal{V},\mathcal{E})$
+5
 1:  for $i$ from $1$ to $M$ do
-2:      Create node $v_{i}\in\mathcal{V}$ using the centroid $p_{i}$ and bounding box size $b_{i}$ of object point cloud $P_{i}$
+12:     Create node $v_{i}\in\mathcal{V}$ using the centroid $p_{i}$ and bounding box size $b_{i}$ of object
+point cloud $P_{i}$
 3:  end for
@@
 6:        $\mathrm{RelsType}_{v}\leftarrow\mathrm{VerticalInContact}(v_{i},v_{j})$
-7:         Add in-contact vertical relationship triplets $(v_{i},v_{j},e_{i,j})$ with $\mathrm{RelsType}_{v}$ to $\mathcal{G}$
+27:        Add in-contact vertical relationship triplets $(v_{i},v_{j},e_{i,j})$ with $\mathrm{RelsType}_{v}$
+to $\mathcal{G}$
 8:     end for
@@
 11:           $\mathrm{RelsType}_{v}\leftarrow\mathrm{VerticalNonContact}(v_{i},v_{k})$
-12:            Add non-contact vertical relationship triplets $(v_{i},v_{k},e_{i,k})$ with $\mathrm{RelsType}_{v}$ to $\mathcal{G}$
+312:           Add non-contact vertical relationship triplets $(v_{i},v_{k},e_{i,k})$ with
+$\mathrm{RelsType}_{v}$ to $\mathcal{G}$
 13:        end for
@@
 16:  for  $v_{i}\in\mathcal{V}$ do
-17:      let ${\{v_{i_{1}},v_{i_{2}},...,v_{i_{N}}\}}$ be the $N$ different nodes with the same in-contact vertical parent node $v_{i}$
+417:     let ${\{v_{i_{1}},v_{i_{2}},...,v_{i_{N}}\}}$ be the $N$ different nodes with the same in-contact
+vertical parent node $v_{i}$
 18:     for $j$ from $1$ to $N$ do
@@
 Algorithm 1 Scene Graph Construction Pipeline
-Figure A.1: Overview of the relationships in SceneVerse. The target object is colored in blue.
-Figure A.2: Examples of object captioning. We color the target object in bold.
-Figure A.3: Examples of scene captioning.
-Figure A.4: Examples of object referral. Note that the green bounding box indicates the target object and yellow bounding box indicates the anchor object(s).
-A.3 Language Generation Details
+0.A.3 Language Generation Details
 In Sec. 3.4, we adopt both templates and LLM to automatically generate scene-language pairs in SceneVerse. More technical details and examples are provided in this section.
-A.3.1 Object Captioning Pipeline
+0.A.3.1 Object Captioning Pipeline
 Object captions aim to provide detailed descriptions of an object’s visual and physical properties, facilitating object-level grounding with its distinctive features.
-The detailed object captioning pipeline is outlined in Alg.
-2.
+The detailed object captioning pipeline is outlined in Algorithm 2.
 Given the multi-view images $\{I_{1},I_{2},\dots,I_{n}\}$, we utilize the point cloud $P_{o}$ of the object $o$ to get the visible points $P_{o,v}^{vis}$ in the images $v$ through rendering.
@@
 In this process, we explicitly instruct the language model to identify and correct the potential errors.
+Figure A.1: Overview of the relationships in SceneVerse. The target object is colored in blue.
+Figure A.2: Examples of object captioning. We color the target object in bold.
+Figure A.3: Examples of scene captioning.
+Figure A.4: Examples of object referral. Note that the green bounding box indicates the target object and yellow bounding box indicates the anchor object(s).
 Input :  $M$ object point clouds $\{P_{1},P_{2},\dots,P_{m}\}$; $N$ multiview images $\{I_{1},I_{2},\dots,I_{n}\}$
@@
 5:        Get the image caption $C_{o,v}$ for $I^{crop}_{o,v}$ using BLIP2 [58]
-6:         Calculate the similarity score $s^{clip}_{o,v}$ between $C_{o,v}$ and $I^{crop}_{o,v}$ with CLIP [66]
+16:        Calculate the similarity score $s^{clip}_{o,v}$ between $C_{o,v}$ and $I^{crop}_{o,v}$ with
+CLIP [78]
 7:        Calculate the occlusion score $s^{occ}_{o,v}=\frac{\#P^{vis}_{o,v}}{\#P_{o}}$
@@
 Algorithm 2 Object Captioning Pipeline
-A.3.2 Automatic Language Generation
-Template-based
-We create diverse templates to generate descriptions for each type of relationship.
+0.A.3.2 Automatic Language Generation
+  1. 1.
+    Template-based We create diverse templates to generate descriptions for each type of relationship.
 We categorized the templates into three types based on the number of objects involved and their spatial relationships.
@@
     Table A.2: Prompts used in SceneVerse.
-Description type
-Prompt
-Object caption
-Summarize caption below.
+    Description type Prompt
+    Object caption Summarize caption below.
 The summary should be a description of the target-object.
@@
     target-object: Bed
-Object referral
-Rewrite the following caption using one random sentence structure. You should give me only one rewritten sentence without explanation.
+    Object referral Rewrite the following caption using one random sentence structure.
+You should give me only one rewritten sentence without explanation.
 caption: The bed is between desk and nightstand.
@@
 caption: The armchair is next to the sofa.
-target-object: Armchair
-anchor-object(s): Sofa
+target-object: Armchair anchor-object(s): Sofa
     Rewrite the following caption using one random sentence structure.
@@
 target-object: Vase
-Scene captioning
-Your task is to provide a summary for a scene from a given scene graph. The scene contains some objects, which compose a scene graph in json format.
+    Scene captioning Your task is to provide a summary for a scene from a given scene graph. The scene contains some objects, which compose a scene graph in json format.
     There are 3 types of descriptions in scene graph: “scene type” denotes the type of the scene.
@@
 You must write using one random sentence structure.
-scene graph: { ’scene_type’: ’Bedroom’, ’object_count’: {’nightstand’:2, ...}, ’relation’: {’nightstand’, ’on’, ’floor’}, {’backback’, ’in front of’, bed}, ...}
-LLM-rephrasing
-To increase description diversity we use the GPT-3.5[61] and Llama[75] for description rephrasing.
+    scene graph: { ‘scene_type’: ‘Bedroom’, ‘object_count’: {‘nightstand’:2, ...}, ‘relation’: {‘nightstand’, ‘on’, ‘floor’}, {‘backback’, ‘in front of’, ‘bed’}, ...}
+  2. 2.
+    LLM-rephrasing To increase description diversity we use the GPT-3.5[73] and Llama[86] for description rephrasing.
 This improves the diversity and naturalness of the template-based descriptions, as is shown in Fig.
@@
 More examples of the scene-language pairs in SceneVerse are shown in Fig. A.2, Fig. A.3 and Fig. A.4.
-Appendix B Model Details
-B.1 Spatial-Attention Transformer
+Appendix 0.B Model Details
+0.B.1 Spatial-Attention Transformer
 In Sec.
@@
 This process could be equivalently interpreted as using the spatial location of objects to adjust the self-attention feature aggregation between objects, making spatially related objects have more attention weights.
-B.2 Pre-training Details
+0.B.2 Pre-training Details
 For training our model GPS, we conduct a two-stage training approach.
@@
 All pre-training experiments are run on 8 NVIDIA-A100 GPUs with the longest pre-training on SceneVerse taking about 2 days.
-Appendix C Experimental Details
+Appendix 0.C Experimental Details
 In this section, we provide details on experimental settings, model implementation, and additional results.
-C.1 3D Visual Grounding
+0.C.1 3D Visual Grounding
 Setting
@@
 For the pre-train setting, we follow the same setting mentioned in Sec.
-B.2.
+0.B.2.
 In the scratch and fine-tuned settings, to fairly compare with other dataset-specific fine-tuned models, we add an additional 2-layer MLP over the object features from the referral grounding self-attention transformer.
 During training, we fine-tune this grounding head together with all model weights for 100 epochs with a learning rate of $1\times 10^{-4}$ for the added projection layer and set all other settings the same as the implementation described in Sec.
-B.2.
-C.2 Zero-shot Transfer
+0.B.2.
+0.C.2 Zero-shot Transfer
 Setting
@@
 This setting serves to test models’ generalization capability in grounding objects with both unseen scenes and unseen texts.
-In the zero-shot text setting, we add generated scene-text pairs in ScanNet and MultiScan into the data used in the zero-shot setting, thereby making the held-out test containing mainly unseen object referrals.
+In the zero-shot text setting, we add generated scene-text pairs in ScanNet and MultiScan into the data used in the zero-shot setting, thereby making the held-out test contain mainly unseen object referrals.
 Implementation
@@
 For the zero-shot setting, we pre-train the model following Sec.
-B.2 without additional grounding heads considering there is no additional training data available in the zero-shot transfer setting.
+0.B.2 without additional grounding heads considering there is no additional training data available in the zero-shot transfer setting.
 In the scratch and zero-shot text setting, we follow the model implementation described in Sec.
-C.1 and add an additional 2-layer MLP over the object features from the self-attention transformer.
+0.C.1 and add an additional 2-layer MLP over the object features from the self-attention transformer.
 We follow the same fine-tuning setting described in Sec.
-C.1.
-Appendix D Additional Results
-In this section, we provide additional experimental results.
-Specifically, we leverage our collected SceneVerse as the pre-training data source for a traditional 3D semantic segmentation task.
-Next, we provide ablative analyses of our model design.
-D.1 Semantic Segmentation
+0.C.1.
+0.C.3 3D question answering
+Setting
+In the 3D-QA experiments, we evaluate all models with only the training sets provided for fine-tuning.
+Following previous works [112], we report model performance on the validation and test sets of ScanQA and the test set of SQA3D.
+Implementation
+As briefly mentioned in Sec.
+5.3, we mainly considered the fine-tuned GPS when comparing with existing methods.
+For all datasets, we initialize our GPS model from a checkpoint pre-trained with 3D visual grounding on SceneVerse.
+We follow 3D-VisTA and add an additional question-answering module over the pre-trained representations for the answer prediction.
+We fine-tune the model for 100 epochs with a learning rate of $1\times 10^{-4}$ for the added question answering head and set all other settings the same as the implementation described in Sec.
+0.B.2.
+0.C.4 Open-vocabulary 3D semantic segmentation
+Setting
+Following RegionPLC [97] proposed by Yang et al., we conduct experiments to assess the performance of SceneVerse on open-vocabulary 3D semantic segmentation (OV-Seg).
+To establish a benchmark for open-vocabulary semantic segmentation, we adopt the experimental setup outlined in PLA [31], as per the methodology of RegionPLC.
+We utilize the annotation-free training setting, as described in RegionPLC, wherein semantic labels for all categories are omitted.
+This approach allows us to evaluate the effectiveness of SceneVerse in facilitating open-vocabulary segmentation without relying on predefined semantic segmentation annotations.
+For evaluation, we compute the mean Intersection over Union (mIoU) and mean accuracy (mAcc) across 17 foreground categories, excluding “wall” and “floor” background classes, as well as the “other furniture” category due to its inherent ambiguity.
+Implementation
+We employ SparseUNet [36] as our 3D backbone network for extracting point features.
+We utilize different variants of SparseUNet, varying the number of channels in the input layer to explore its impact on performance.
+For text feature extraction, we employ CLIP [40] text encoder.
+To align the extracted features from the 3D scene encoder and the text encoder, we incorporate a vision-language adapter.
+The only supervision comes from the point-discriminative contrastive loss proposed by RegionPLC.
+During training, we employ the AdamW optimizer to update model parameters.
+We train the model from scratch for 500 epochs, utilizing a learning rate of $1\times 10^{-3}$.
+Additionally, we incorporate a warm-up period of 200 steps and a cosine annealing scheme for learning rate scheduling, with a minimum learning rate ratio of $1\times 10^{-5}$.
+Appendix 0.D Additional Results
+0.D.1 Semantic Segmentation
 Setting
@@
 Table A.3: Semantic segmentation results on ScanNet validation set.
-normal-$\dagger$ denotes model trained with surface normals as an additional input.
+$\dagger$ denotes model trained with surface normals as an additional input.
 S3D indicates models initialized with the original Swin3D model weights pre-trained on Structured3D provided by Yang et al.
@@
 We believe these results serve as strong pieces of evidence validating the effectiveness of data scaling in SceneVerse and also its potential benefit for all 3D tasks in addition to 3D visual grounding.
-D.1.1 Model Ablation
-In this section, we provide ablative analyses of our multi-level contrastive alignment design.
-We mainly consider removing objectives in our model as ablations.
-We choose the referral-object-level alignment objective as the default setting and consider removing: (i) object-level alignment objective, (ii) masked language modeling objective, and (iii) scene-level alignment objective.
-For removing the object-level alignment objective, we remove the first stage pre-training of the object point cloud encoder and jointly learn this module within the referral-object-level alignment.
-As shown in Tab.
-A.4, we test different model settings on the SceneVerse-val without additional fine-tuning.
-First, we show that the scene-level alignment objective is crucial for referral object grounding in SceneVerse-val with the $\sim$5performance drop.
-Similar observations could be made for the model trained without object-level alignment ($\sim$2drop) and masked language modeling objective ($\sim$1.5drop).
-We believe these results affirm the effectiveness of our overall model design.
-Table A.4: Model ablation of our model on SceneVerse-val.
-Obj-lvl MLM Scene-lvl Overall Easy Hard
-64.8 75.4 48.7
-65.2 77.1 47.4
-62.4 73.4 45.8
-66.9 77.8 50.3
+0.D.2 Qualitative Results
+We provide the qualitative results of 3D vision-language grounding in Fig. A.5 and the results of open-vocabulary semantic segmentation in Fig. A.6.
+Figure A.5: Qualitative results of GPS on 3D visual-language grounding. We visualize the incorrect predictions in red and the correct predictions or ground truths in green.
+Figure A.6: Qualitative results of open-vocabulary 3D semantic segmentation (OV-Seg). We label the object class in ScanNet-20’s vocabulary in blue, and unseen class in ScanNet-20 in green.